### v2

# Cross-Validation with Antithetic Gaussian Randomization

Abstract
We introduce a new cross-validation method based on an equicorrelated Gaussian randomization scheme. The method is well-suited for problems where sample splitting is infeasible, such as when data violate the assumption of independent and identical distribution. Even when sample splitting is possible, our method offers a computationally efficient alternative for estimating the prediction error, achieving comparable or even lower error than standard cross-validation in a few train-test repetitions.
Drawing inspiration from recent techniques like data-fission and data-thinning, our method constructs train-test data pairs using externally generated Gaussian randomization variables. The key innovation lies in a carefully designed correlation structure among the randomization variables, which we refer to as antithetic Gaussian randomization. In theory, we show that this correlation is crucial in ensuring that the variance of our estimator remains bounded while allowing the bias to vanish. Through simulations on various data types and loss functions, we highlight the advantages of our antithetic Gaussian randomization scheme over both independent randomization and standard cross-validation, where the bias-variance tradeoff depends heavily on the number of folds.
1 Introduction
Estimating prediction error is a fundamental task in statistics and machine learning, essential for assessing how well a model generalizes to unseen data, selecting tuning parameters during estimation, and comparing different models. Cross-validation is one of the most widely used tools for this purpose. In its standard form, the data is partitioned into independent subsamples or “folds” and prediction error is obtained by averaging the empirical errors from the test folds. The popularity of cross-validation is easy to understand—it is versatile and applies to a wide range of loss functions and data types, due to its assumption-light nature.
The standard form of cross-validation is, however, not suitable for all types of data, especially when the assumptions of independent and identically distributed observations are not satisfied. For example, in regression settings with influential observations, a subset of samples may fail to adequately represent the full dataset. When dealing with categorical response variables or covariates, sample splitting may lead to imbalanced folds, potentially omitting rare categories from some folds entirely. For time series or spatially correlated data, splitting the data can disrupt the inherent temporal or spatial structure. In such cases, standard cross-validated estimators of prediction error can be misleading and can result in unreliable models for downstream tasks.
In this paper, we address this issue by introducing a novel cross-validation method that eliminates the need for sample splitting. Instead, the train-test folds in our method are created with externally generated Gaussian randomization variables. The method is governed by two user-specified parameters, $\alpha$ and $K$. The first parameter, $\alpha\in\mathbb{R}^{+}$, is akin to the proportion of held-out samples in standard cross-validation. The second parameter, $K\in\mathbb{N}$, specifies the number of train-test repetitions over which estimates of prediction error are averaged.
The proposed method is as follows: we generate $K$ randomization variables from an equicorrelated and degenerate normal distribution with a zero-sum constraint. By adding a $\sqrt{\alpha}$-scaled version of these randomization variables to the sufficient statistics, we create $K$ train-test data pairs. Prediction error is then estimated using these pairs in a manner similar to standard cross-validation. For example, consider normal data $Y\in\mathbb{R}^{n}$ with a covariance matrix $\sigma^{2}I_{n}$. In this case, the train-test data for the $k$-th repetition are constructed as
$Y_{\text{train}}^{(k)}=Y+\sqrt{\alpha}\omega^{(k)},\quad Y_{\text{test}}^{(k)}=Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)},$ (1)
where $\omega^{(k)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$, for $k\in[K]=\{1,2,\ldots,K\}$, are equicorrelated Gaussian randomization variables that sum to zero. In this paper, we extend this approach to handle a wide range of loss functions and data types, as long as the sufficient statistics for the unknown parameters in the loss function are asymptotically normal.
1.1 Highlights of our method
The performance of any cross-validation method, measured by mean squared error (MSE), depends on the bias-variance tradeoff, which is influenced by both the proportion of held-out data during training and the number of train-test repetitions. In standard cross-validation, this tradeoff is controlled by the number of folds. Our cross-validation method is particularly appealing because it provides two distinct levers to control the bias and variance of the associated estimator for prediction error. This is outlined below:
  1. 1.
    Direct control of bias via $\bm{\alpha}$: The parameter $\alpha$ controls the bias introduced by estimating the prediction function on noisier training data, with the bias decaying to 00 as $\alpha$ decreases. Unlike standard cross-validation, where bias is controlled by the number of folds, the parameter $\alpha$ in our method is independent of the number of train-test repetitions, $K$. This separation provides a significant advantage: by averaging empirical estimates of prediction error over just $K$ train-test repetitions—where $K$ can be as few as two—our method, with a small $\alpha$, can achieve a bias comparable to that of leave-one-out (LOO) cross-validation. Thus, even when sample splitting is feasible, the new cross-validated estimator offers a computationally efficient alternative for estimating prediction error.
  2. 2.
    Stable variance for finite $\mathbf{K}$: A key strength of the proposed estimator, as supported by our theoretical analysis, is its stable variance for any finite $K$, even as the bias decays to zero with decreasing $\alpha$. This contrasts with standard cross-validation, where reducing bias often results in increased variance. The stability of the variance is due to the carefully designed correlation structure of the external Gaussian randomization variables. Following the literature on variance reduction techniques for Monte Carlo methods, e.g., Craiu and Meng, (2005), we view our randomization approach as an “extreme antithesis”, where the correlation between any pair of randomization variables takes the most negative value possible.
To the best of our knowledge, this work is the first to investigate the potential of an antithetic Gaussian randomization approach for cross-validation. It provides a unique and a computationally efficient solution for reducing bias in the estimation of prediction errors, while maintaining a stable variance. Figure 1 showcases the performance of our new cross-validated estimator by comparing its mean squared error (MSE) against that of standard cross-validation estimators. In this example, we focus on estimating the prediction error for an isotonic regression problem. Our method uses only two train-test repetitions ($K=2$) with $\alpha=0.01$, while classic cross-validation is performed with $K=2$ folds and $K=100$ folds, the latter corresponding to leave-one-out (LOO) cross-validation. Remarkably, our estimator achieves a smaller MSE than LOO cross-validation while being $50$ times more computationally efficient. More details about this example, along with extensive numerical results that examine the effects of $\alpha$ and $K$, are presented in Section 6.
Figure 1: Mean squared error (MSE) for estimating prediction error in an isotonic regression problem using a simulated dataset. From left to right, the methods shown are classic 2-fold CV, LOO CV, and the proposed method with $K=2$ and $\alpha=0.01$. Additional details are provided in Section 6.
1.2 Related work and contributions
Our cross-validation proposal is inspired by several recently introduced randomized methods that provide alternatives to traditional sample splitting for tasks such as model validation, selective inference, and risk estimation. These alternatives include data-fission and data-thinning techniques by Rasines and Young, (2023); Leiner et al., (2023); Neufeld et al., (2024); Dharamshi et al., (2024), methods employing Gaussian randomization for selective inference tasks, as considered in Dai et al., (2023); Tian and Taylor, (2018); Panigrahi and Taylor, (2022); Huang et al., (2023), and randomized methods by Oliveira et al., (2021, 2022); Fry and Taylor, (2023) for unbiased estimation of risk and prediction errors.
Our cross-validation method, like data fission or data thinning techniques, is naturally suited for problems where sample splitting is infeasible. However, unlike these existing methods, which use different randomization schemes tailored to specific parametric distributions, our approach employs the same Gaussian randomization scheme for different loss functions and justifies their use within a relatively assumption-light framework. In fact, the idea of employing alternative forms of randomization for cross-validation is by no means new. For example, Brown et al., (2013) described a “nonstandard cross-validation method” for the Gaussian sequence model. They propose using a single train-test split of the form (1) for estimation and hyperparameter tuning. This construction is closely related to our proposal when we only use two “folds” and it is also a key motivating example of data fission (Leiner et al.,, 2023). Similarly, the multifold thinning approach in Neufeld et al., (2024) proposed the use of correlated Gaussian randomization variables for cross-validation in the normal means problem. However, their correlation structure differs from the antithetic randomization scheme proposed in our work, a distinction that we highlight in our concluding discussion.
Similar randomization schemes, where Gaussian noise is added to the sufficient statistic, have been prominent in the selective inference literature. For example, in the randomized lasso estimators by Panigrahi and Taylor, (2022); Panigrahi et al., (2024) and the randomized group lasso estimators by Panigrahi et al., (2023), Gaussian noise is added to the objective function of the optimization problem. This randomized scheme is indeed equivalent to adding normal variables to the sufficient statistic in Gaussian regression models. The randomization framework for generalized linear models (GLMs) developed by Liu and Panigrahi, (2023) for selective inference with distributed data employs the same antithetic approach as presented in this paper, though it serves a different purpose. As a natural by-product, our proposal here can also be seen to offer a way to perform cross-validation in these randomized problems, particularly for selecting optimal tuning parameters that determine the amount of sparsity in the selected model.
Among the methods reviewed, the one most closely related to our work is the coupled bootstrap (CB) estimator proposed by Oliveira et al., (2021) for normal data, which we discuss in detail in the next section. The CB estimator computes prediction error using randomized train-test data constructed with independent Gaussian randomization variables. A key advantage of our cross-validated estimator over the CB estimator lies in its substantial variance reduction, achieved by deliberately using an antithetic Gaussian randomization scheme.
Here is a summary of our main contributions in the remainder of the paper:
  1. 1.
    In Section 2, we review the CB estimator for the normal means problem with a quadratic loss function and introduce our cross-validated estimator, based on antithetic Gaussian randomization variables.
  2. 2.
    In Section 3, we analyze the mean squared error of the proposed estimator as $\alpha$, the parameter controlling bias, approaches zero. Our theory demonstrates that we can obtain unbiased estimates of prediction error as $\alpha\to 0$, while ensuring that the variance of our estimator remains stable even with vanishingly small $\alpha$. In contrast to the CB estimator, which requires increasing $K$ as $\alpha$ decreases, our method can achieve the same variance with significantly smaller $K$. This analysis highlights the benefits of employing a carefully chosen antithetic randomization scheme instead of an independent randomization scheme.
  3. 3.
    In Section 4, we establish connections between the proposed estimator and classical risk estimators, such as Stein’s Unbiased Risk Estimator (SURE) and its variants for exponential families. Notably, our estimator can be viewed as replacing the divergence term in SURE by the divergence of a Gaussian-smoothed version of the prediction function.
  4. 4.
    In Section 5, we extend our cross-validation framework to accommodate more general loss functions, including those commonly used in fitting GLMs, such as logistic regression. Under the assumption that the sufficient statistics are asymptotically normal and satisfy certain regularity conditions, we demonstrate that the mean squared error analysis generalizes to a broader class of loss functions.
  5. 5.
    In Section 6, we provide simulation results comparing our proposed framework to standard cross-validation, the coupled bootstrap, and SURE. The proposed method performs effectively across various data types, loss functions, and prediction algorithms. It eliminates the need for sample splitting, manual tuning of the bias-variance tradeoff, or differentiating the prediction function. Additionally, the method is computationally efficient, requiring us to conduct only a small number of train-test repetitions.
  6. 6.
    In Section 7, we conclude with a discussion of potential extensions and new directions for the proposed method.
2 Basic setup and the proposed estimator
Here, we outline the setup of our problem. We assume that the response vector $Y=(Y_{1},\ldots,Y_{n})^{\intercal}\in\mathbb{R}^{n}$ is drawn from a distribution ${\mathbb{P}}_{n}$, while the predictors or covariates are treated as fixed. A prediction function $g$ is trained on this data. Given a loss function ${\mathcal{L}}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, our goal is to evaluate the performance of this prediction function on unseen test data ${\widetilde{Y}}$, which is an independent copy of the observed data $Y$. Our estimand of interest is the expected prediction error, defined as
$\mathrm{PE}(g)=\mathbb{E}\left[{\mathcal{L}}(g(Y),{\widetilde{Y}})\right],$
where the expectation is taken over both the training data $Y$ and the testing data ${\widetilde{Y}}$.
The most common approach to estimating prediction error involves splitting the sample space. In this approach, the $n$ observations $(Y_{1},\ldots,Y_{n})$ are randomly divided into two non-overlapping subsets, $Y^{(1)}$ and $Y^{(2)}$. The prediction function $g$ is trained on the first subset $Y^{(1)}$, and its performance is evaluated on the second subset $Y^{(2)}$, resulting in the prediction error estimator
${\mathcal{L}}\left(g(Y^{(1)}),Y^{(2)}\right).$ (2)
A more data-efficient approach to the same problem employs the $K$-fold cross-validation (CV), where the $n$ observations are randomly partitioned into $K$ non-overlapping folds, denoted by $Y^{(k)}$ for $k\in[K]$. Each fold is used for both training and testing, and the prediction error is finally estimated as
$\frac{1}{K}\sum_{k=1}^{K}{\mathcal{L}}(g(Y^{(-k)}),Y^{(k)}).$
Here, $Y^{(-k)}$, the complement of the $k$-th fold $Y^{(k)}$, is used for training the prediction function $g$, and the held-out fold, $Y^{(k)}$, is used for evaluating the predictive performance of $g$ in the $k$-th repetition. The bias-variance tradeoff in standard cross-validation depends on the number of folds $K$, and practitioners often face the challenge of selecting the optimal value of $K$ to achieve an effective tradeoff between the bias and variance of the resulting estimator.
This paper introduces a novel approach to cross-validation that constructs train-test data using external randomization variables. Unlike standard cross-validation, our method addresses the bias-variance tradeoff by controlling two separate parameters: $\alpha$, which controls bias, and $K$, which controls variance. The advantages of this new form of cross-validation, with two user-specified parameters, will become evident through our analysis of the mean squared error.
Before presenting our method, we first review the coupled bootstrap (CB) estimator proposed by Oliveira et al., (2021), which also utilizes external randomization variables to construct train-test data.
2.1 Review of coupled bootstrap (CB)
The CB estimator (Oliveira et al.,, 2021) aims to estimate the risk in the normal means problem, where the response vector $Y\in\mathbb{R}^{n}$ is assumed to follow the normal distribution $\mathcal{N}(\theta,\sigma^{2}I_{n})$, with a known variance $\sigma^{2}$. In this work, we focus on the prediction error for a prediction function $g$, defined as
$\mathrm{PE}(g)=\mathbb{E}\left[\|g(Y)-{\widetilde{Y}}\|_{2}^{2}\right],$ (3)
where ${\widetilde{Y}}\sim\mathcal{N}(\theta,\sigma^{2}I_{n})$ is an independent copy of $Y$. Note that our estimand differs from the risk by a constant in the normal means problem.
To estimate $\mathrm{PE}(g)$, the CB method generates $K$ independent Gaussian randomization variables
$\widetilde{\omega}^{(1)},\widetilde{\omega}^{(2)},\ldots,\widetilde{\omega}^{(K)}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,\sigma^{2}I_{n}).$
For each $k\in[K]$ and a parameter $\alpha\in\mathbb{R}^{+}$, two randomized copies of $Y$ are constructed as
$\widetilde{Y}^{(k)}_{\text{train}}=Y+\sqrt{\alpha}\widetilde{\omega}^{(k)},\quad\widetilde{Y}^{(k)}_{\text{test}}=Y-\dfrac{1}{\sqrt{\alpha}}\widetilde{\omega}^{(k)},$ (4)
where, by construction, the two vectors are distributed as
$\begin{pmatrix}\widetilde{Y}^{(k)}_{\text{train}}\\ \widetilde{Y}^{(k)}_{\text{test}}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\theta\\ \theta\end{pmatrix},\begin{bmatrix}\sigma^{2}(1+\alpha)I_{n}&0_{n,n}\\ 0_{n,n}&\sigma^{2}(1+\alpha^{-1})I_{n})\end{bmatrix}\right).$
The prediction error based on the $k$-th train-test pair is computed as
${\text{CB}}_{\alpha}^{(k)}=\|\widetilde{Y}^{(k)}_{\text{test}}-g(\widetilde{Y}^{(k)}_{\text{train}})\|_{2}^{2}-\frac{1}{\alpha}\|\widetilde{\omega}^{(k)}\|_{2}^{2},$ (5)
where the second term, $\|\widetilde{\omega}^{(k)}\|_{2}^{2}/\alpha$, adjusts for the difference between the variance of the randomized test data and the variance of the original data $Y$. Finally, the CB estimator is obtained by averaging over $K$ independent draws of the Gaussian randomization variables
${\text{CB}}_{\alpha}=\frac{1}{K}\sum_{k=1}^{K}{\text{CB}}_{\alpha}^{(k)}.$
Since ${\widetilde{Y}}^{(k)}_{\text{train}}\sim\mathcal{N}(\theta,(1+\alpha)\sigma^{2}I_{n})$, straightforward calculations show that the CB estimator is unbiased for a noise-inflated version of the prediction error
$\mathrm{PE}_{\alpha}(g)=\mathbb{E}\left[\|g(Y)-{\widetilde{Y}}\|_{2}^{2}\right],\text{ where }Y\sim\mathcal{N}(\theta,(1+\alpha)\sigma^{2}I_{n}),\;{\widetilde{Y}}\sim\mathcal{N}(\theta,\sigma^{2}I_{n}).$
This estimand corresponds to the prediction error when $g$ is trained on noisier data, with variance inflated by a factor of $(1+\alpha)$. The estimator $\mathrm{CB}_{\alpha}$ is, therefore, biased for the true prediction error $\mathrm{PE}(g)$, defined in Equation (3). However, the bias—the difference between the noise-inflated prediction error $\mathrm{PE}_{\alpha}(g)$ and the original estimand $\mathrm{PE}(g)$—converges to zero as the parameter $\alpha$ approaches zero. Nevertheless, as in standard train-test splitting, a bias-variance tradeoff arises here: reducing the bias by decreasing $\alpha$ comes at the expense of increased variance. As shown in Oliveira et al., (2021), the variance of the CB estimator is of order $O((K\alpha)^{-1})$. This implies that, for any finite $K$, the variance of the CB estimator becomes unbounded as the bias decreases to 00.
We address this limitation of the CB estimator by introducing a randomization scheme with a carefully chosen correlation structure, which we refer to as an “antithetic” randomization scheme.
2.2 Antithetic randomization
In our antithetic randomization scheme, we generate $K$ ($K>1$) randomization variables as follows:
$\omega^{(1)},\ldots,\omega^{(K)}\sim\mathcal{N}(0,\sigma^{2}I_{n}),\text{ where }\text{Cov}(\omega^{(j)},\omega^{(k)})=-\frac{\sigma^{2}}{K-1}I_{n}\text{ for }j\neq k.$ (6)
We make two important observations about this distribution. First, the normal distribution in (6) is degenerate. This is because the variance of the sum of the randomization variables is zero, i.e., $\text{Var}\left(\sum_{k=1}^{K}\omega^{(k)}\right)=0$. Combined with fact that the randomization variables have zero mean, this imposes the following zero-sum constraint on these randomization variables:
$\sum_{k=1}^{K}\omega^{(k)}=0.$ (7)
Second, for a $K$-by-$K$ correlation matrix where all off-diagonal entries are equal, the range of possible correlation is
$[-\frac{1}{K-1},1].$
Therefore, our randomization scheme takes the most negative correlation possible, which is why we refer to it as “antithetic”.
For a fixed $\alpha\in\mathbb{R}^{+}$, we construct randomized train-test copies of the data $Y$ as
$\begin{pmatrix}Y^{(k)}_{\text{train}}\\ Y^{(k)}_{\text{test}}\end{pmatrix}=\begin{pmatrix}Y-\sqrt{\alpha}\displaystyle\sum_{j\neq k}\omega^{(k)}\\ Y-\dfrac{1}{\sqrt{\alpha}}\omega^{(k)}\end{pmatrix}=\begin{pmatrix}Y+\sqrt{\alpha}\omega^{(k)}\\ Y-\dfrac{1}{\sqrt{\alpha}}\omega^{(k)}\end{pmatrix},\;\text{ for }k\in[K],$
where the second equality is due to the zero-sum constraint in (7). This approach mimics the standard $K$-fold cross-validation in that, when pooling the train (or test) data from all $K$ folds, the randomization variables cancel out, thereby recovering the original data $Y$.
Our cross-validated estimator $\mathrm{CV}_{\alpha}$ is then defined as
${\text{CV}}_{\alpha}=\frac{1}{K}\sum_{k=1}^{K}{\text{CV}}_{\alpha}^{(k)},$ (8)
where
${\text{CV}}_{\alpha}^{(k)}$ $=\|Y^{(k)}_{\text{test}}-g(Y^{(k)}_{\text{train}})\|_{2}^{2}-\frac{1}{\alpha}\|\omega^{(k)}\|_{2}^{2}.$
The key distinction between the CB estimator and the proposed estimator lies in the randomization scheme. In the coupled bootstrap method, the randomization variables $\widetilde{\omega}^{(1)},\ldots,\widetilde{\omega}^{(K)}$ are independent. In contrast, our method employs correlated randomization variables $\omega^{(1)},\ldots,\omega^{(K)}$. As will be shown in the next section, this correlation leads to a significant variance reduction, ensuring that the variance of our cross-validated estimator remains bounded as $\alpha\to 0$, at which point the bias of our estimator also vanishes.
3 Mean squared error analysis
In this section, we analyze the mean squared error (MSE) of the proposed estimator $\mathrm{CV}_{\alpha}$ (8) for estimating the prediction error $\mathrm{PE}(g)$ (3) in the normal means problem. The MSE can be decomposed into bias and variance as
$\mathbb{E}\left[(\mathrm{CV}_{\alpha}-\mathrm{PE}(g))^{2}\right]$ =$=\left\{\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]-\mathrm{PE}(g)\right\}^{2}+\operatorname{Var}\left[\mathrm{CV}_{\alpha}\right]$
=$=\left\{\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]-\mathrm{PE}(g)\right\}^{2}+\mathbb{E}\left[\operatorname{Var}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right]+\operatorname{Var}\left[\mathbb{E}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right].$ (9)
We study the bias $\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]-\mathrm{PE}(g)$ in Section 3.1, and the reducible variance $\mathbb{E}\left[\operatorname{Var}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right]$ and irreducible variance $\operatorname{Var}\left[\mathbb{E}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right]$ in Section 3.2.
3.1 Bias
We show that the bias $\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]-\mathrm{PE}(g)$ can be made arbitrarily small as $\alpha$ approaches zero, under the mild condition that $\|g(Y)\|_{2}^{2}$ is integrable. This result follows directly from the “approximation to the identity” property of the Gaussian density, as stated in Lemma 3.1 below.
Let $\varphi_{\sigma^{2}}$ denote the density of the normal distribution $\mathcal{N}(0,\sigma^{2}I_{n})$. Let $f*\varphi_{\sigma^{2}}$ denote the convolution of an integrable function $f$ with $\varphi_{\sigma^{2}}$, which is defined as
$f*\varphi_{\sigma^{2}}(y):=\int f(y-z)\varphi_{\sigma^{2}}(z)\mathrm{d}z.$
Lemma 3.1 (Approximation to the identity).
Let $f$ be an integrable function under the Gaussian distribution $\mathcal{N}(\theta,\sigma^{2}I_{n})$. Then
$f*\varphi_{\alpha\sigma^{2}}(Y)\stackrel{{\scriptstyle L_{1}}}{{\to}}f(Y)\text{ as }\alpha\to 0.$
Proof of Lemma 3.1.
This is a direct application of Lemma D.1 and Lemma D.2 in the Appendix. ∎
Lemma 3.1 states that the convolution of a function with $\varphi_{\alpha\sigma^{2}}$ is close to the original function in the $L_{1}$ sense as $\alpha\to 0$. In the context of our problem, this lemma implies that
$\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}g(Y)$
as $\alpha\to 0$, which is the key to showing that the bias of our estimator converges to 0 as $\alpha$ approaches zero. The result is formalized in the following theorem.
Theorem 3.2 (Bias).
Assume that $\mathbb{E}\left[\|g(Y)\|_{2}^{2}\right]<\infty$. Then we have
$\lim_{\alpha\to 0}\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]=\mathrm{PE}(g).$
Proof of Theorem 3.2.
Since $\mathbb{E}\left[\mathrm{CV}_{\alpha}\right]=\mathbb{E}\left[\mathrm{CV}_{\alpha}^{(k)}\right]$, it is sufficient to compute the expectation of $\mathrm{CV}_{\alpha}^{(k)}$. Observe that
$\mathbb{E}\left[\mathrm{CV}_{\alpha}^{(k)}\right]$ $=\mathbb{E}\left[\|Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)}-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}-\frac{\|\omega^{(k)}\|_{2}^{2}}{\alpha}\right]$
$=\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}-2(Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)})^{\intercal}g(Y+\sqrt{\alpha}\omega^{(k)})\right]$
$\ \ \ \ +\mathbb{E}\left[\|Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)}\|_{2}^{2}\right]-\mathbb{E}\left[\frac{\|\omega^{(k)}\|_{2}^{2}}{\alpha}\right]$
$=\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}\right]-2\mathbb{E}\left[(Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)})\right]^{\intercal}\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega^{(k)})\right]+\mathbb{E}\left[\|Y\|_{2}^{2}\right]$
$=\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}\right]-2\mathbb{E}\left[Y\right]^{\intercal}\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega^{(k)})\right]+\mathbb{E}\left[\|Y\|_{2}^{2}\right]$
[…]
Note that
$\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega^{(k})\|_{2}^{2}\mid Y\right]=\|g\|_{2}^{2}*\varphi_{\alpha\sigma^{2}}(Y),$
which converges in $L_{1}$ to $\|g(Y)\|_{2}^{2}$ as $\alpha\to 0$, by Lemma 3.1. Similarly, applying Lemma 3.1 to the function $g_{i}(Y)$ for $1\leq i\leq n$ shows that that $\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega^{(k)})\mid Y\right]$ converges in $L_{1}$ to $g(Y)$.
This establishes that, as $\alpha\to 0$,
$\mathbb{E}\left[\mathrm{CV}_{\alpha}^{(k)}\right]\to\mathbb{E}\left[\|g(Y)\|_{2}^{2}\right]-2\mathbb{E}\left[Y\right]^{\intercal}\mathbb{E}\left[g(Y)\right]+\mathbb{E}\left[\|Y\|_{2}^{2}\right].$
The right-hand-side equals $\mathrm{PE}(g)=\mathbb{E}\left[\|\widetilde{Y}-g(Y)\|_{2}^{2}\right]$, where $\widetilde{Y}$ is an independent copy of $Y$. This completes the proof. ∎
Consequently, the proposed estimator $\mathrm{CV}_{\alpha}$ has vanishingly small bias when $\alpha$ is chosen to be small. In standard $K$-fold cross-validation, reducing bias typically requires increasing $K$, which leads to higher computational costs and often greater variance. In contrast, our estimator achieves low bias by simply using a small $\alpha$, without the need to increase $K$. More importantly, as we will demonstrate next, unlike the coupled bootstrap method, decreasing $\alpha$ does not increase the variance of our estimator.
3.2 Variance reduction with antithetic randomization
To analyze the variance of the proposed estimator $\mathrm{CV}_{\alpha}$, we impose a mild smoothness condition on the prediction function $g$. This condition is the weak differentiability assumption considered in the classical SURE estimator (Stein,, 1981).
Assumption 1 (Weak differentiability).
All components $g_{i}$ ($1\leq i\leq n$) of $g$ are weakly differentiable. That is, there exists a function $\nabla g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, the weak derivative of $g_{i}$, such that
$g_{i}(y+z)-g_{i}(y)=\int_{0}^{1}z\cdot\nabla g_{i}(y+tz)\mathrm{d}t,$
for almost all $y,z\in\mathbb{R}^{n}$. Denote the Jacobian matrix of $g$ as $\nabla g\in\mathbb{R}^{n\times n}$, where the $i$-th row is equal to $\nabla g_{i}$.
This class of functions encompasses many well-known estimators, including the ridge estimator, the lasso estimator, the group lasso estimator, and the generalized lasso estimator; see, for example, the paper by Tibshirani and Taylor, (2012).
The following theorem provides the expression for the reducible variance of $\mathrm{CV}_{\alpha}$ as $\alpha$ approaches zero.
Theorem 3.3 (Reducible variance).
Suppose that Assumption 1 holds. Furthermore, let $\mathbb{E}\left[\|g(Y)\|_{2}^{4}\right]<\infty$, $\mathbb{E}\left[\|\nabla g(Y)\|_{F}^{2}\right]<\infty$. Then, we have that
$\lim_{\alpha\to 0}\mathbb{E}\left[\operatorname{Var}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right]=\frac{4\sigma^{4}}{K-1}\mathbb{E}\left[\|\nabla g(Y)\|_{F}^{2}+\operatorname{tr}(\nabla g(Y)^{2})\right].$
Remark 3.1.
Theorem 3.3 implies that the reducible variance of our cross-validated estimator remains bounded for any fixed $K>1$, even as $\alpha\to 0$. In contrast, the CB estimator, based on independent randomization variables, has a reducible variance of order $O(\frac{1}{K\alpha})$, which diverges to $\infty$ as $\alpha\to 0$ for any finite $K$.
We provide a sketch of the proof here to illustrate the role of antithetic randomization in achieving this reduction in variance, with the detailed proof deferred to Appendix A.1.
Proof sketch of Theorem 3.3.
We first write
$\mathrm{CV}_{\alpha}$ $=\frac{1}{K}\sum_{k=1}^{K}\|Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)}-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}-\frac{1}{\alpha}\|\omega^{(k)}\|_{2}^{2}$
$=\underbrace{\frac{1}{K}\sum_{k=1}^{K}\|Y-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}}_{(I)}+\underbrace{\frac{1}{K}\sum_{k=1}^{K}\frac{2}{\sqrt{\alpha}}\langle\omega^{(k)},g(Y+\sqrt{\alpha}\omega^{(k)})\rangle}_{(II)}$ (10)
$\qquad\qquad-\underbrace{\frac{2}{K}\sum_{k=1}^{K}\langle Y,\frac{1}{\sqrt{\alpha}}\omega^{(k)}\rangle}_{=0}.$
Note that the last term is 0 because of the zero-sum property of the antithetic randomization variables, i.e., $\sum_{k=1}^{K}\omega^{(k)}=0$.
Note that
$\operatorname{Var}\left[\mathrm{CV}_{\alpha}\mid Y\right]=\operatorname{Var}\left[(I)\mid Y\right]+\operatorname{Var}\left[(II)\mid Y\right]+2\mathrm{Cov}[{(I),(II)\mid Y}].$
For the first summation $(I)$, we show that
$\operatorname{Var}\left[(I)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}0.$
This is because we can write this conditional variance as the convolution of an integrable function with the Gaussian density $\varphi_{\alpha\sigma^{2}}$, which converges in $L_{1}$ to 0, by the “approximation to identity property of the Gaussian density”, as stated in Lemma 3.1.
For the second summation $(II)$, we have by the definition of weak differentiability that
$(II)$ $=\frac{2}{K\sqrt{\alpha}}\sum_{k=1}^{K}\langle\omega^{(k)},g(Y)+\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(k)})^{\intercal}(\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\rangle$
$=\frac{2}{K}\sum_{k=1}^{K}{\omega^{(k)}}^{\intercal}\left[\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\right]\omega^{(k)}.$ (11)
The last equality is due to the fact that $\sum_{k=1}^{K}\omega^{(k)}=0$, which forces the term
$\frac{2}{K\sqrt{\alpha}}\sum_{k=1}^{K}\langle\omega^{(k)},g(Y)\rangle$
term to vanish.
The “approximation to identity property” is applied again to show that
$\operatorname{Var}\left[(II)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}\operatorname{Var}\left[\frac{2}{K}\sum_{k=1}^{K}{\omega^{(k)}}^{\intercal}\nabla g(Y)\omega^{(k)}\mid Y\right].$
The right-hand-side in the last display is the variance of a quadratic form of the Gaussian vector $(\omega^{(1)},\ldots,\omega^{(K)})$, which has a closed form as given in the statement of the Theorem.
Lastly, $\mathrm{Cov}[{(I),(II)\mid Y}]\stackrel{{\scriptstyle L_{1}}}{{\to}}0$ by noting that
$\mathbb{E}\left[\mathrm{Cov}[{(I),(II)\mid Y}]\right]$ $\leq\mathbb{E}\left[\sqrt{\operatorname{Var}\left[(I)\mid Y\right]\operatorname{Var}\left[(II)\mid Y\right]}\right]$
$\leq\sqrt{\mathbb{E}\left[\operatorname{Var}\left[(I)\mid Y\right]\right]}\sqrt{\mathbb{E}\left[\operatorname{Var}\left[(II)\mid Y\right]\right]}.$
The first inequality in the above display follows by applying the Cauchy-Schwarz inequality
$\mathrm{Cov}[{(I),(II)\mid Y}]\leq\sqrt{\operatorname{Var}\left[(I)\mid Y\right]\operatorname{Var}\left[(II)\mid Y\right]}.$
Finally, to complete the analysis of variance of our estimator, we provide the limit of the irreducible variance.
Theorem 3.4 (Irreducible variance).
Under the same assumptions as in Theorem 3.3, we have that
$\lim_{\alpha\to 0}\operatorname{Var}\left[\mathbb{E}\left[\mathrm{CV}_{\alpha}\mid Y\right]\right]=\operatorname{Var}\left[\|Y-g(Y)\|_{2}^{2}+2\sigma^{2}\operatorname{tr}(\nabla g(Y))\right].$
The proof is provided in Appendix A.2.
Combining the bias-variance results in Theorem 3.2, 3.3 and 3.4, we find that, as $\alpha\to 0$,
$\text{MSE}(\mathrm{CV}_{\alpha})\to\operatorname{Var}\left[\|Y-g(Y)\|_{2}^{2}+2\sigma^{2}\operatorname{tr}(\nabla g(Y))\right]+\frac{4\sigma^{4}}{K-1}\mathbb{E}\left[\|\nabla g(Y)\|_{F}^{2}+\operatorname{tr}(\nabla g(Y)^{2})\right].$
Recall that the MSE of the CB estimator is dominated by a term of order $O(1/\alpha)$ as $\alpha\to 0$ for any finite $K$. In contrast, the MSE of the proposed estimator remains bounded, leading to the following corollary.
Corollary 1.
Under the same assumptions as in Theorem 3.3, for any finite $K>1$, we have that
$\lim_{\alpha\to 0}\left\{\mathrm{MSE}(\mathrm{CV}_{\alpha})-\mathrm{MSE}(\mathrm{CB}_{\alpha})\right\}=-\infty.$
This result indicates that our cross-validated estimator offers an infinite efficiency gain over the coupled bootstrap method. Moreover, by selecting a small $\alpha$, we can make the bias arbitrarily small while ensuring that the variance does not blow up. This stability in variance underscores the advantages of the proposed antithetic randomization scheme.
4 Connection with SURE
For the normal means problem, a well-known method for risk estimation is Stein’s Unbiased Risk Estimator (SURE) (Stein,, 1981), which is defined as
$\mathrm{SURE}(g)=\|Y-g(Y)\|_{2}^{2}+2\sigma^{2}\nabla\cdot g(Y),$
where the divergence of $g$ is given by $\nabla\cdot g(Y)=\operatorname{tr}(\nabla g(Y))$. SURE is commonly used to estimate the quadratic risk $\mathbb{E}\left[\|\theta-g(Y)\|_{2}^{2}\right]$. In the normal means problem, the quadratic risk and the prediction error differ only by a constant $n\sigma^{2}$. Therefore, we analyze SURE here as an estimator of the prediction error $\mathrm{PE}(g)$.
Under Assumption 1, along with the conditions that $\mathbb{E}\left[\|g(Y)\|_{2}^{2}\right]<\infty$ and $\mathbb{E}\left[|\nabla_{i}g_{i}(Y)|\right]<\infty$, the SURE estimator is unbiased for the prediction error $\mathrm{PE}(g)$. The unbiased-ness of SURE follows directly from Stein’s identity for Gaussian distributions:
$\mathbb{E}\left[(Y-\theta)^{\intercal}g(Y)\right]=\sigma^{2}\mathbb{E}\left[\nabla\cdot g(Y)\right].$
We argue that our estimator $\mathrm{CV}_{\alpha}$ closely resembles SURE, despite being motivated from a completely different perspective. Recall from Equation (10) that our estimator can be expressed as
$\mathrm{CV}_{\alpha}=\frac{1}{K}\sum_{k=1}^{K}\|Y-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}+\frac{1}{K}\sum_{k=1}^{K}\frac{2}{\sqrt{\alpha}}(\omega^{(k)})^{\intercal}g(Y+\sqrt{\alpha}\omega^{(k)}).$ (12)
For small $\alpha$, we claim that
$\mathbb{E}\left[\mathrm{CV}_{\alpha}\mid Y\right]\approx\|Y-g(Y)\|_{2}^{2}+2\sigma^{2}\nabla\cdot g(Y)=\mathrm{SURE}(g).$
This is due to the following reasons. By Lemma 3.1, the conditional expectation of the first term in (12), $\mathbb{E}\left[\|Y-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}\mid Y\right]$, converges in $L_{1}$ as $\alpha\to 0$ to $\|Y-g(Y)\|_{2}^{2}$, which is the first term in $\text{SURE}(g)$. Moreover, according to Equation (11), the second term in (12) equals
$\frac{1}{K}\sum_{k=1}^{K}\frac{2}{\sqrt{\alpha}}(\omega^{(k)})^{\intercal}g(Y+\sqrt{\alpha}\omega^{(k)})$ $=\frac{2}{K}\sum_{k=1}^{K}{\omega^{(k)}}^{\intercal}\left[\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\right]\omega^{(k)},$
By a reasoning similar to Lemma 3.1, we can show that as $\alpha\to 0$
$\mathbb{E}\left[\frac{2}{K}\sum_{k=1}^{K}{\omega^{(k)}}^{\intercal}\left[\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\right]\omega^{(k)}\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}2\sigma^{2}\nabla\cdot g(Y),$
which corresponds to the second term in $\text{SURE}(g)$. Consequently, after integrating out the randomization variables, the proposed estimator $\mathrm{CV}_{\alpha}$ converges to SURE$(g)$ in $L_{1}$ as $\alpha\to 0$.
Furthermore, even for a positive $\alpha$, the proposed estimator remains closely related to SURE. In fact, we argue that the proposed estimator corresponds to the SURE applied to a convolution-smoothed version of the prediction function $g$. To see this, consider the expression for $\mathrm{CV}_{\alpha}$ in Equation (12), and replace the term $g(Y+\sqrt{\alpha}\omega^{(k)})$ with its conditional expectation $\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega)\mid Y\right]$, where the expectation is over $\omega\sim\mathcal{N}(0,\sigma^{2}I_{n})$. This leads to the noise-free version of our estimator:
$\overline{\mathrm{CV}}_{\alpha}=\|Y-\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega)\mid Y\right]\|_{2}^{2}+\frac{2}{\sqrt{\alpha}}\mathbb{E}\left[\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right],$ (13)
In other words, $\overline{\mathrm{CV}}_{\alpha}$ corresponds to $\mathrm{CV}_{\alpha}$ with the randomness from $\omega^{(k)}$’s marginalized out.
The following result states that the noise-free version $\overline{\mathrm{CV}}_{\alpha}$ of the proposed estimator, coincides with the SURE when $g$ is replaced by its convolution-smoothed version $g*\varphi_{\alpha\sigma^{2}}$.
Proposition 4.1 (Connection with SURE).
It holds that
$\overline{\mathrm{CV}}_{\alpha}=\mathrm{SURE}(g*\varphi_{\alpha\sigma^{2}}).$ (14)
The proof is provided in Appendix B.1. Two remarks are in order.
Remark 4.1.
When SURE is applicable, the proposed estimator behaves similarly to SURE when $\alpha$ is small. Our estimator, however, does not require computing the divergence term $\nabla\cdot g$, which may not be available in closed form for many estimators. This makes $\mathrm{CV}_{\alpha}$ a more practical choice in such scenarios.
Remark 4.2.
When SURE is not applicable, such as when $g$ is not weakly differentiable, the proposed estimator $\mathrm{CV}_{\alpha}$ remains well-defined. In these cases, $\mathrm{CV}_{\alpha}$ behaves as though applying SURE to the infinitely differentiable, convolution-smoothed estimator $g*\varphi_{\alpha\sigma^{2}}$. This connection with SURE provides further justification for the proposed method, providing a solution in settings where SURE is not applicable.
4.1 Generalization to exponential families
Given the connection between $\mathrm{CV}_{\alpha}$ and SURE, we can naturally generalize our estimator to other exponential families, using the more general version of Stein’s identity for this larger family of distributions.
Suppose $Y\in\mathbb{R}^{n}$ follows the exponential family distribution with density
$p(Y)=\exp(\theta^{\intercal}Y-A(\theta))\cdot h(Y),$
where $\theta\in\mathbb{R}^{n}$ is the natural parameter, $A(\theta)$ is the log-partition function, and $h$ is the base measure. Let $g(Y)$ be an estimator of $\theta$. Our goal is to estimate the risk under the quadratic loss $\mathbb{E}\left[\|\theta-g(Y)\|_{2}^{2}\right]$. Since $\|\theta\|_{2}^{2}$ is a constant not depending on the estimator and $\mathbb{E}\left[\|g(Y)\|_{2}^{2}\right]$ can be estimated by $\|g(Y)\|_{2}^{2}$, the task reduces to estimating the cross term $\mathbb{E}\left[\theta^{\intercal}g(Y)\right]$.
Stein’s identity (see, for example, Eldar, (2008)):
$\mathbb{E}\left[\theta^{\intercal}g(Y)\right]=-\mathbb{E}\left[\nabla\cdot g(Y)+g(Y)^{\intercal}\nabla\log h(Y)\right]$ (15)
implies that
$-\nabla\cdot g(Y)-g(Y)^{\intercal}\nabla\log h(Y)$
is an unbiased estimator of $\mathbb{E}\left[\theta^{\intercal}g(Y)\right]$. However, this estimator involves the divergence term $\nabla\cdot g(Y)$, which is often unavailable.
In line with our earlier arguments, we propose to approximate the divergence term $\nabla\cdot g$ by its convolution-smoothed version $\nabla\cdot(g*\varphi_{\alpha\sigma^{2}})$. This term can then be estimated using the Monte Carlo estimator
$\frac{1}{K\sqrt{\alpha}}\sum_{k=1}^{K}{\omega^{(k)}}^{\intercal}g(y+\sqrt{\alpha}\omega^{(k)}),$
where
$\omega^{(k)}\sim\mathcal{N}(0,I_{n}),\ \operatorname{Cov}\left[\omega^{(j)},\omega^{(k)}\right]=-\frac{1}{K-1}I_{n}\text{ for }j\neq k.$
The advantages of using antithetic randomization extend here as well, ensuring that the variance remains bounded even as $\alpha\to 0$, at which point the bias also vanishes.
5 Extensions beyond the quadratic loss
In this section, we extend our cross-validation method to handle more general loss functions, where the sufficient statistic in the loss function is asymptotically normal. To emphasize the dependence on the sample size, we add subscripts $n$ to the data, the estimand, and the estimator. Later in the section, we analyze the bias and variance of the proposed estimator in the asymptotic regime as $n\to\infty$.
Suppose the data $Y=Y_{n}$ is generated from an exponential family with density:
$p_{n}(Y_{n}\mid\theta_{n})=\exp\left\{\sqrt{n}(\theta_{n}^{\intercal}S_{n}(Y_{n})-A_{n}(\theta_{n}))\right\}\cdot h_{n}(Y_{n}),$
where $\theta_{n}$ is the $p$-dimensional natural parameter. Note, in this formulation, the sufficient statistic $S_{n}=S_{n}(Y_{n})$ and the log-partition function $A_{n}(\theta_{n})$ are scaled by $1/\sqrt{n}$. We consider a loss function derived from the negative log-likelihood of this density, which is given by
${\mathcal{L}}(\theta_{n},Y_{n})=A_{n}(\theta_{n})-\theta_{n}^{\intercal}S_{n}(Y_{n})-\frac{1}{\sqrt{n}}\log h_{n}(Y_{n}).$ (16)
This setup accommodates the loss functions typically used in fitting generalized linear models (GLM). Throughout this section, we assume the existence of a sequence of $p\times p$ positive definite matrix $H_{n}$ and $\mu_{n}\in\mathbb{R}^{p}$ such that
$H_{n}^{-1/2}(S_{n}-\mu_{n})\stackrel{{\scriptstyle d}}{{\Rightarrow}}\mathcal{N}(0,I_{p}).$ (17)
The asymptotic normality assumption holds in GLMs under regularity conditions as established in Fahrmeir and Kaufmann, (1985).
5.1 Cross-validated estimator
Suppose that $g(S_{n})$ is an estimator of $\theta_{n}$, which depends on the data only through the sufficient statistic $S_{n}$. As before, we define the prediction error as the expectation of the loss function:
$\mathrm{PE}_{n}(g)=\mathbb{E}\left[{\mathcal{L}}(g(S_{n}),\widetilde{Y}_{n})\right]=\mathbb{E}\left[A_{n}(g(S_{n}))-g(S_{n})^{\intercal}\widetilde{S}_{n}-n^{-1/2}\log h_{n}({\widetilde{Y}}_{n})\right],$
where $\widetilde{Y}_{n}$ is an independent copy of $Y$, and $\widetilde{S}_{n}=S_{n}(\widetilde{Y}_{n})$ is the sufficient statistic of $\widetilde{Y}_{n}$.
We define the rescaled sufficient statistics as
$T_{n}=H_{n}^{-1/2}S_{n},\quad\widetilde{T}_{n}=H_{n}^{-1/2}\widetilde{S}_{n}.$
By Equation (17), the asymptotic distributions of $T_{n}-H_{n}^{-1/2}\mu_{n}$ and $\widetilde{T}_{n}-H_{n}^{-1/2}\mu_{n}$ are $\mathcal{N}(0,I_{p})$. Let
$\mathfrak{g}_{n}(T_{n})=(H_{n}^{1/2})^{\intercal}g(H_{n}^{1/2}T_{n}),\quad\mathfrak{A}_{n}(T_{n})=A_{n}(g(H_{n}^{1/2}T_{n})),$
such that
$A_{n}(g(S_{n}))=\mathfrak{A}_{n}(T_{n}),\quad g(S_{n})^{\intercal}\widetilde{S}_{n}=\mathfrak{g}(T_{n})^{\intercal}\widetilde{T}_{n}.$
With these notations, we can rewrite the prediction error as
$\mathrm{PE}_{n}(g)=\mathbb{E}\left[\mathfrak{A}_{n}(T_{n})-\mathfrak{g}_{n}(T_{n})^{\intercal}\widetilde{T}_{n}\right]-\mathbb{E}\left[n^{-1/2}\log h_{n}(Y_{n})\right].$ (18)
The second expectation in our estimand, $\mathbb{E}\left[n^{-1/2}\log h_{n}(Y_{n})\right]$, can be easily estimated by $n^{-1/2}\log h_{n}(Y_{n})$. The first expectation is taken over $T_{n}$ and $\widetilde{T}_{n}$, which are asymptotically normal with identity covariance. Thus, the problem reduces to a form analogous to the normal means example discussed earlier, except that $T_{n}$ is not exactly normal but asymptotically normal.
We apply the same idea as before, constructing the train-test pair of randomized data as
$T_{n}+\sqrt{\alpha}\omega\quad\text{and}\quad T_{n}-\frac{1}{\sqrt{\alpha}}\omega,\quad\text{where }\omega\sim\mathcal{N}(0,I_{p}),$
for $\alpha\in\mathbb{R}^{+}$. Clearly, the train-test data are asymptotically independent. We train the prediction function on $T_{n}+\sqrt{\alpha}\omega$ and evaluate its performance on $T_{n}-\frac{1}{\sqrt{\alpha}}\omega$, leading to the following estimate of $\mathrm{PE}_{n}(g)$:
${\mathfrak{A}}_{n}(T_{n}+\sqrt{\alpha}\omega)-{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)^{\intercal}(T_{n}-\frac{1}{\sqrt{\alpha}}\omega)-n^{-1/2}\log h_{n}(Y_{n}).$
We propose to repeat this procedure $K>1$ times, with randomization variables $\omega^{(1)},\ldots,\omega^{(K)}$ generated using the antithetic scheme described in (6), i.e.,
$\omega^{(k)}\sim\mathcal{N}(0,I_{p}),\quad\operatorname{Cov}\left[\omega^{(j)},\omega^{(k)}\right]=\frac{-1}{K-1}I_{p}\,\text{ for }j\neq k.$ (19)
Averaging over the $K$ draws of randomization, we obtain the cross-validated estimator
$\mathrm{CV}_{n,\alpha}=\frac{1}{K}\sum_{k=1}^{K}$$\Big{\{}\mathfrak{A}_{n}(T_{n}+\sqrt{\alpha}\omega^{(k)})-\mathfrak{g}_{n}(T_{n}+\sqrt{\alpha}\omega^{(k)})^{\intercal}(T_{n}-\frac{1}{\sqrt{\alpha}}\omega^{(k)})\Big{\}}$ (20)
$\quad-n^{-1/2}\log h_{n}(Y_{n}).$
Note that we could equivalently work with the sufficient statistics $S_{n}$ directly, without rescaling them to $T_{n}$. In this case, the randomization variables would be introduced with a marginal covariance matrix equal to $H_{n}$, while maintaining the same antithetic correlation structure used throughout our method. With the rescaling of the sufficient statistics, we instead work with randomization variables that have an identity covariance matrix, which simplifies the presentation.
As we demonstrate next, the proposed estimator exhibits similar desirable bias-variance properties to those in the normal means problem. Specifically, the asymptotic bias vanishes as $\alpha\to 0$, and the variance remains bounded, which is again a consequence of the antithetic randomization scheme.
5.2 Mean squared error analysis
To conduct the mean squared error analysis of our cross-validated estimator $\mathrm{CV}_{n,\alpha}$, we require some additional assumptions on the sufficient statistics $T_{n}$.
For a weakly differentiable $\mathbb{R}^{p}$-valued function $g$ and a $p$-dimensional vector $\mu$, define
$({\mathcal{T}}_{\mu}g)(x)=\langle g(x),\mu-x\rangle+\nabla\cdot g(x).$
For a normal random variable $X\sim\mathcal{N}(\mu,I_{p})$, it follows that $\mathbb{E}\left[({\mathcal{T}}_{\mu}g)(X)\right]=0$, which recovers Stein’s identity.
Let $\mathbb{Q}_{n}$ represent the distribution of the rescaled sufficient statistics, $T_{n}$, with density $q_{n}$ and expectation $m_{n}=H_{n}^{-1/2}\mu_{n}$.
Assumption 2.
Assume that
$\lim_{n\to\infty}\mathbb{E}\left[({\mathcal{T}}_{m_{n}}g_{n})(T_{n})\right]=0$
where
$({\mathcal{T}}_{m_{n}}g)(x)=\langle g(x),m_{n}-x\rangle+\nabla\cdot g(x).$
Under a distribution $\mathbb{Q}_{n}$ that is not normal, note that $\mathbb{E}\left[({\mathcal{T}}_{m_{n}}g_{n})(T_{n})\right]$ is no longer exactly zero. This quantity, known as Stein’s measure of non-normality, forms the basis for the notion of Stein’s discrepancy; see, for example, the paper by Gorham and Mackey, (2015). Assumption 2 requires that the sufficient statistics exhibit vanishingly small Stein’s discrepancy as $n$ goes to infinity. For example, given that the sufficient statistics are asymptotically normal, this condition holds if $\|T_{n}\|_{q}^{q}$ is also uniformly integrable, and both functions $\langle g(x),x\rangle$, $\nabla\cdot g(x)$ grow slower than $\|x\|_{q}^{q}$ for some $q>0$.
Assumption 3.
Assume that there exist constants $N_{0}>0$ and $C>0$ such that, for all $n\geq N_{0}$, the density $q_{n}$ of $T_{n}$ satisfies
$|\log q_{n}(x)-\log q_{n}(x^{\prime})|\leq C\|x-x^{\prime}\|_{2}^{2}.$
The condition in Assumption 3 is automatically satisfied if the density of the sufficient statistics converges to a normal density.
Now we are ready to show that the bias and variance results established in Section 3 for exactly normal data carry over to our estimator based on asymptotically normal sufficient statistics. In particular, we show that the asymptotic bias is 0 as $\alpha\to 0$ and $n\to\infty$. Moreover, the variance remains bounded as $\alpha\to 0$.
Theorem 5.1 (Bias).
Let Assumptions 1, 2, and 3 hold. In addition, assume that $\mathbb{E}\left[|{\mathfrak{A}}_{n}(T_{n})|\right]<\infty$, $\mathbb{E}\left[\|{\mathfrak{g}}_{n}(T_{n})\|_{2}^{2}\right]<\infty$, and $\mathbb{E}\left[|\nabla{\mathfrak{g}}_{n}(T_{n})|\right]<\infty$. Then
$\lim_{n\to\infty}\lim_{\alpha\to 0}\Big{|}\mathbb{E}\left[\mathrm{CV}_{n,\alpha}\right]-\mathrm{PE}_{n}(g)\Big{|}=0.$
Theorem 5.2 (Reducible variance).
Let Assumptions 1 and 3 hold. In addition, assume that $\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n})^{2}\right]<\infty$, $\mathbb{E}\left[\|{\mathfrak{g}}_{n}(T_{n})\|_{2}^{4}\right]<\infty$, and $\mathbb{E}\left[\|\nabla{\mathfrak{g}}_{n}(T_{n})\|_{F}^{2}\right]<\infty$. When $n\geq N_{0}$, we have
$\lim_{\alpha\to 0}\mathbb{E}\left[\operatorname{Var}\left[\mathrm{CV}_{n,\alpha}\mid Y_{n}\right]\right]=\frac{1}{K-1}\mathbb{E}\left[\|{\mathfrak{g}}_{n}(T_{n})\|_{F}^{2}+\operatorname{tr}(\nabla{\mathfrak{g}}_{n}(T_{n})^{2})\right].$
The proofs are provided in Appendix C.1 and C.2.
We conclude this section by presenting the widely used logistic regression for classification as an example within our setup. Simulations in the next section illustrate the performance of our cross-validated estimator in this example.
Example 5.3 (Logistic regression).
In logistic regression, the negative log-likelihood is given by
$\sum_{i=1}^{n}-y_{i}x_{i}^{\intercal}\theta+\log(1+e^{x_{i}^{\intercal}\theta}),$
where the data follows the model
$y_{i}\sim\mathrm{Bernoulli}(\pi(x_{i}^{\intercal}\theta_{n})),$
where $\pi(x)=\frac{1}{1+\exp(-x)}$ is the sigmoid function. In the above-described setup, we scale the log-likelihood by $1/\sqrt{n}$, resulting in the loss function
${\mathcal{L}}(\theta_{n},Y_{n})=-\theta_{n}^{\intercal}S_{n}+A_{n}(\theta_{n}),$
where
$S_{n}=\frac{1}{\sqrt{n}}X_{n}^{\intercal}Y_{n},\quad A_{n}(\theta_{n})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\log(1+e^{x_{i}^{\top}\theta_{n}}).$
Under mild regularity conditions, such as those outlined in Fahrmeir and Kaufmann, (1985), it can be shown that the sufficient statistic $S_{n}$ follows an asymptotically normal distribution. The covariance matrix in this limiting normal distribution is equal to $H_{n}=\dfrac{1}{n}X_{n}^{\intercal}W_{n}X_{n}$, where $W_{n}$ is a diagonal matrix with entries $\pi(x_{i}^{\intercal}\theta_{n})(1-\pi(x_{i}^{\intercal}\theta_{n}))$. Clearly, this covariance depends on the true underlying parameter. However, our method performs effectively as long as we can estimate this matrix, as demonstrated in the simulations presented in Section 6.
6 Numerical experiments
In this section, we evaluate our method empirically on simulated data and compare our method with competing estimators for estimating prediction errors.
6.1 Effects of $\alpha$ and $K$
In our first example, we examine the effects of $\alpha$ and $K$, the two parameters that control the bias and variance of the proposed cross-validation estimator.
We consider an isotonic regression problem with $n=100$ observations, in which the one-dimensional predictors are drawn independently from $\textnormal{Uniform}(0,1)$ and are treated as fixed throughout this experiment. The responses are then generated independently as $Y_{i}\sim\mathcal{N}(f^{*}(X_{i}),\sigma^{2})$, where the true function $f^{*}(x)=2\cdot\lceil 5x\rceil-6$ and the noise level $\sigma^{2}=1$. Visualizations of the function $f^{*}$ and simulated data points are provided in Figure 2. We apply isotonic regression to the data to estimate the true prediction function $f^{*}(x)$. A snapshot of this example was previously presented in the introduction to illustrate comparisons between our method and the standard CV method. Throughout this section, we will refer to our method as “Antithetic CV”.
Figure 2: In the isotonic regression simulations, the data are generated based on the function $f^{*}$, shown in solid line. An example of a simulated dataset is displayed as scatter points.
We compare the MSE of the proposed Antithetic CV method and the Coupled Bootstrap (CB) method across various values of $\alpha$ and $K$. In the left panel of Figure 3, $\alpha$ is fixed at 0.1, and $K$ is varied from 2 to 32. In the right panel, $K$ is fixed at 8, and $\alpha$ is varied from 0.005 to 0.5. In all comparisons, standard CV uses the same value of $K$ for train-test sample splitting, as the other two methods.
From this figure, we make a few key observations:
  1. 1.
    The antithetic CV method consistently outperforms standard $K$-fold CV and the coupled bootstrap method across all configurations of $(\alpha,K)$.
  2. 2.
    In the left panel of Figure 3, the MSE of all methods decreases for all three methods as the number of train-test repetitions $K$ increases. For the coupled bootstrap and antithetic CV methods, this reduction can be attributed to the decrease in reducible variance with increasing $K$.
  3. 3.
    In the right panel of Figure 3, we observe that the MSE of CB decreases as $\alpha$ increases. This is because the MSE of CB is dominated by the variance when $\alpha$ is small. As a result, even though the bias of CB decreases with smaller $\alpha$, the increase in variance outweighs this benefit, leading to a higher overall MSE. In contrast, for antithetic CV, we note that the MSE either decreases or does not increase as $\alpha$ becomes smaller. This is expected as our approach achieves bias reduction without paying a price in terms of the variance.
Figure 3: Mean squared error of standard $K$-fold CV, coupled bootstrap with $K$ repetitions, and the proposed Antithetic CV methods with $K$ repetitions. Left panel: MSE with $\alpha=0.1$ and varying $K$. Right panel: MSE with $K=8$ and varying $\alpha$.
6.2 Logistic regression
We illustrate the effectiveness of the proposed method for estimating prediction errors in GLMs, as described in Section 5.
We consider a logistic regression example with both continuous and categorical predictors and $n=100$. The $4$ continuous features $X^{\text{cts}}$ are generated from the standard Gaussian distribution independently, and each of the $2$ categorical features, $X^{\text{cat}}$ has 3 classes drawn with probability $[0.1,0.1,0.8]$ independently. Given the predictors, the response $Y_{i}$ is drawn from a Bernoulli distribution with mean parameter $(1+e^{-\eta_{i}})^{-1}$, where
$\eta_{i}=\beta^{\intercal}X_{i}^{\text{cts}}+\sum_{j=1}^{2}\sum_{k=1}^{3}\gamma_{jk}\mathbf{1}\{X_{ij}^{\text{cat}}=k\},$
with $\beta=(1,-1,1,-1)$ and $\gamma_{j}=(1/2,-1/2,0)$ for $j=1,2$. The regression coefficients are estimated by fitting a logistic regression model, where we have used one-hot encoding for the categorical features. The goal is to estimate the prediction error, which in this case is the expectation of the negative log likelihood of the logistic regression model.
The proposed antithetic CV approach, like thinning methods for Gaussian data, assumes knowledge of the covariance matrix of the noise. Here, we demonstrate empirically that the proposed method continues to perform quite well even when we use a plug-in estimate for the covariance matrix. Specifically, we estimate the unknown covariance matrix of the sufficient statistic $H_{n}$ by $\widehat{H}_{n}=\frac{1}{n}X^{\intercal}\widehat{W}X$, where $\widehat{W}$ is the diagonal matrix whose $i$-th diagonal entry is $\pi(x_{i}^{\intercal}\widehat{\theta})(1-\pi(x_{i}^{\intercal}\widehat{\theta}))$ and where $\widehat{\theta}$ is the logistic regression solution. We then use the antithetic CV method introduced in Section 5 with $H_{n}$ replaced by $\widehat{H}_{n}$.
We also consider an alternative method that substitutes the antithetic Gaussian variables in our approach with independent Gaussian variables. This method mirrors the CB method in the normal means problem, and is referred to as “independent randomization”. We set $\alpha=0.1$ for both the antithetic CV method and the method based on a Gaussian randomization scheme with independent variables. In addition, similar to our previous comparisons, we also apply the standard $K$-fold CV estimator in this example. We consider two commonly used values, $K=10$ and $K=20$, which are standard choices in practice.
The MSE of all three estimators is presented in Figure 4. The proposed antithetic CV method achieves a much smaller MSE compared to the standard $K$-fold CV. Furthermore, we observe again a clear advantage of antithetic Gaussian randomization over the independent randomization scheme. These results highlight the substantial improvement in the quality of prediction error estimation that can be achieved through the carefully chosen correlated randomization scheme.
Figure 4: MSE in estimating the prediction error in the logistic regression example. We set $\alpha=0.1$ and consider $K=10$ and 20.
6.3 Multi-layer perceptron regression
In this example, we apply the antithetic CV method to estimate the prediction error of a black-box neural network. The responses are drawn independently as $y_{i}\sim\mathcal{N}(f^{*}(x_{i}),\sigma^{2})$, where $f^{*}$ is the Friedman #1 function (Friedman,, 1991), defined as:
$f^{*}(x)=10\sin(\pi x_{1}x_{2})+20(x_{3}-\frac{1}{2})^{2}+10x_{4}+5x_{5},$
and $\sigma^{2}=1$. We generate $10$-dimensional predictors $x_{i}$ independently and uniformly from $[0,1]^{10}$. We fix $n=1000$. A multi-layer perceptron (MLP) prediction model is fitted using the Python package scikit-learn (Pedregosa et al.,, 2011). The MLP consists of two hidden layers, each with 64 units. The training configuration includes a maximum of $2000$ epochs, a batch size of $64$, and an $L_{2}$ regularization parameter of $0.01$, with all other parameters set to their default values.
We evaluate the proposed antithetic CV method alongside the coupled bootstrap (CB) method described in Section 2, with $\alpha=0.1$. As done in the previous instance of logistic regression, we report results for $K=10$ and $K=20$. Consistent with the previous results, from Figure 5, the proposed cross-validation method achieves the smallest MSE when compared against classic $K$-fold CV and the CB method. While the performance of CB improves with a larger $K$, this comes at a much higher computational cost. To elaborate further, hyperparameter tuning is a key step in nearly all machine learning methods and is typically performed using grid search. However, as the number of hyperparameters increases, this process can quickly become computationally prohibitive. In such cases, we believe that our antithetic cross-validation method provides a more computationally efficient alternative to standard cross-validation, making it particularly well-suited for hyperparameter tuning in resource-constrained environments.
Figure 5: MSE in estimating the prediction error in the MLP regression example. We consider $K=10$ and 20. For CB and antithetic CV, $\alpha$ is set to be 0.1.
7 Discussion
We conclude by discussing related research directions and potential applications of the antithetic randomization scheme in our work.
Multi-fold data thinning
As noted in the introduction, a related concept called multi-fold data thinning (MDT) was proposed in Neufeld et al., (2024). For a vector $Y\sim\mathcal{N}(\theta,\sigma^{2}I_{n})$, the MDT method generates $X^{1},\ldots,X^{K}$ such that $\sum_{k=1}^{K}X^{k}=Y$, $\mathbb{E}\left[X^{(k)}\mid Y\right]=\varepsilon_{k}Y$, $\operatorname{Var}\left[X^{(k)}\mid Y\right]=\varepsilon_{k}(1-\varepsilon_{k})\sigma^{2}$, and $\operatorname{Cov}\left[X^{(j)},X^{(k)}\mid Y\right]=-\varepsilon_{j}\varepsilon_{k}\sigma^{2}$, where $\varepsilon_{k}>0$ and $\sum_{k=1}^{K}\varepsilon_{k}=1$. This method treats $Y^{(k)}_{\mathrm{MDT}}=Y-X^{(k)}$ as the training data and $X^{(k)}$ as the testing data.
If we take $\varepsilon_{k}=1/K$ for all $k$, then the training data $Y^{(k)}_{\mathrm{MDT}}$ follow a joint Gaussian distribution with:
$\mathbb{E}\left[Y^{(k)}_{\mathrm{MDT}}\right]=\frac{K-1}{K}\theta,\quad\operatorname{Cov}\left[Y^{(j)}_{\mathrm{MDT}},Y^{(k)}_{\mathrm{MDT}}\right]=\left(1-\frac{2}{K}+\frac{\delta_{j,k}}{K}\right)\sigma^{2}I_{n}.$
Instead, the training data in the $k$-th fold of our cross-validation method equals $Y_{\mathrm{train}}^{(k)}=Y+\sqrt{\alpha}\omega^{(k)}$, where the joint distribution of the training data across the folds is equal to:
$\mathbb{E}\left[Y_{\mathrm{train}}^{(k)}\right]=\theta,\quad\operatorname{Cov}\left[Y_{\mathrm{train}}^{(j)},Y_{\mathrm{train}}^{(k)}\right]=\left(1-\frac{\alpha}{K-1}+\frac{\alpha K}{K-1}\delta_{j,k}\right)\sigma^{2}I_{n}.$
Even if we multiply $Y^{(k)}_{\mathrm{MDT}}$ by $\frac{K}{K-1}$ to match the mean, the covariance
$\operatorname{Cov}\left[\frac{K}{K-1}Y^{(j)}_{\mathrm{MDT}},\,\frac{K}{K-1}Y^{(k)}_{\mathrm{MDT}}\right]=\left(\frac{K(K-2)}{(K-1)^{2}}+\frac{K}{(K-1)^{2}}\delta_{j,k}\right)\sigma^{2}I_{n}$
is different from that in our method. More importantly, our proposed method introduces an additional parameter, $\alpha$, enabling direct control over the bias of the estimator. In contrast, the MDT method relies solely on a single parameter, $K$, to balance the bias-variance tradeoff, like the standard cross-validation. Moreover, Neufeld et al., (2024) did not provide a quantitative analysis of the bias and variance of the MDT estimator.
Our paper focuses on problems where the sufficient statistics are either normal or asymptotically normal. As future work, it would be interesting to investigate whether similar antithetic randomization schemes could be extended to other convolution-closed distributions, as discussed in the aforementioned reference.
Simultaneous perturbation
The randomization approach proposed in this paper is also connected to simultaneous perturbation methods commonly employed in black-box optimization. Suppose that the objective function is $F(\theta)$ for $\theta\in\mathbb{R}^{d}$, whose gradient is not available. Consider its Gaussian-smoothed version $\mathbb{E}_{\omega\sim\mathcal{N}(0,I)}\left[F(\theta+\sigma\omega)\right]$, whose gradient takes the form
$\nabla_{\theta}\mathbb{E}_{\omega\sim\mathcal{N}(0,I)}\left[F(\theta+\sigma\omega)\right]=\frac{1}{\sigma}\mathbb{E}_{\omega\sim\mathcal{N}(0,I_{d})}\left[F(\theta+\sigma\omega)\omega\right].$
This expectation can be then approximated by sampling $\omega^{(k)}\sim\mathcal{N}(0,I_{d})$ and computing
$\widehat{g}(\theta)=\frac{1}{K\sigma}\sum_{k=1}^{K}F(\theta+\sigma\omega^{(k)})\omega^{(k)}.$
The estimated gradient $\widehat{g}(\theta)$ can be used in any gradient-based optimization algorithm. This method is known as simultaneous perturbation stochastic approximation (Spall,, 1992) or random gradient-free optimization (Nesterov and Spokoiny,, 2017), and it is widely used in black-box optimization tasks such as reinforcement learning (Salimans et al.,, 2017) and fine tuning large language models (Malladi et al.,, 2023).
In these methods, the random perturbations $\omega^{(k)}$ are sampled independently of one another. It is known that the bias of the estimator vanishes as $\sigma\to 0$, while the variance increases as $\sigma$ decreases. The antithetic randomization approach introduced in this paper could be directly applied to gradient estimation in this context, potentially improving the performance of downstream optimization tasks. We leave this as a potential direction for future work.
Appendix A Proofs for Section 3
A.1 Proof of Theorem 3.3
Proof of Theorem 3.3.
We first write
$\mathrm{CV}_{\alpha}$ $=\frac{1}{K}\sum_{k=1}^{K}\|Y-\frac{1}{\sqrt{\alpha}}\omega^{(k)}-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}-\frac{1}{\alpha}\|\omega^{(k)}\|_{2}^{2}$
$=\underbrace{\frac{1}{K}\sum_{k=1}^{K}\left[\|Y-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}\right]}_{(I)}+\underbrace{\frac{1}{K}\sum_{k=1}^{K}\frac{2}{\sqrt{\alpha}}\langle\omega^{(k)},g(Y+\sqrt{\alpha}\omega^{(k)})\rangle}_{(II)}.$
By Lemma A.1, $\operatorname{Var}\left[(I)\mid y\right]$ converges in $L_{1}$ to 0. By Lemma A.2, $\operatorname{Var}\left[(II)\mid Y\right]$ converges in $L_{1}$ to $\operatorname{Var}\left[\frac{2}{K}\sum_{k=1}^{K}(\omega^{(k)})^{\intercal}\nabla g(Y)\omega^{(k)}\mid Y\right]$. When $j\neq k$, $\operatorname{Cov}\left[\omega^{(j)},\omega^{(k)}\right]=\rho\sigma^{2}I$ where $\rho=-\frac{1}{K-1}$. So we have
$\operatorname{Var}\left[\frac{1}{K}\sum_{k}(\omega^{(k)})^{\intercal}\nabla g(Y)\omega^{(k)}\mid Y\right]$
$\qquad=\frac{1}{K^{2}}\left(K\cdot\operatorname{Var}\left[\omega^{\intercal}\nabla g(Y)\omega\right]+K(K-1)\operatorname{Cov}\left[(\omega^{(1)})^{\intercal}\nabla g(Y)\omega^{(1)},(\omega^{(2)})^{\intercal}\nabla g(Y)\omega^{(2)}\right]\right).$
By Lemma D.5,
$\operatorname{Var}\left[\omega^{\intercal}\nabla g(Y)\omega\right]=\sigma^{4}(\|\nabla g(Y)\|_{F}^{2}+\operatorname{tr}(\nabla g(Y)^{2})),$
$\operatorname{Cov}\left[(\omega^{(1)})^{\intercal}\nabla g(Y)\omega^{(1)},(\omega^{(2)})^{\intercal}\nabla g(Y)\omega^{(2)}\right]=\frac{1}{(K-1)^{2}}\operatorname{Var}\left[\omega^{\intercal}\nabla g(Y)\omega\right].$
Therefore,
$\operatorname{Var}\left[\frac{1}{K}\sum_{k}(\omega^{(k)})^{\intercal}\nabla g(Y)\omega^{(k)}\mid Y\right]$ $=\frac{1}{K^{2}}\left(K+K(K-1)\frac{1}{(K-1)^{2}}\right)\operatorname{Var}\left[\omega^{\intercal}\nabla g(Y)\omega\right]$
$=\frac{\sigma^{4}}{K-1}(\|\nabla g(Y)\|_{F}^{2}+\operatorname{tr}(\nabla g(Y)^{2})).$
This completes the proof. ∎
Lemma A.1 (First term $(I)$).
Assume that $\mathbb{E}\left[\|g(Y)\|_{2}^{4}\right]<\infty$. Then as $\alpha\to 0$,
$\operatorname{Var}\left[\frac{1}{K}\sum_{k=1}^{K}\|Y-g(Y+\sqrt{\alpha}\omega^{(k)})\|_{2}^{2}\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}0.$
Proof.
Because $\operatorname{Var}\left[X_{1}+\ldots+X_{K}\right]\leq K\cdot(\operatorname{Var}\left[X_{1}\right]+\ldots+\operatorname{Var}\left[X_{K}\right])$, it suffices to show that as $\alpha\to 0$,
$\operatorname{Var}\left[\|Y-g(Y+\sqrt{\alpha}\omega)^{\intercal}\|_{2}^{2}\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}0,$
where $\omega\sim\mathcal{N}(0,\sigma^{2}I_{n})$. It suffices to show that $\mathbb{E}\left[\|Y-g(Y+\sqrt{\alpha}\omega)\|_{2}^{4}\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}\|Y-g(Y)\|_{2}^{4}$, which is equivalent to showing that $\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega)\|_{2}^{4}\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}\|g(Y)\|_{2}^{4}$. This is true by applying Lemma 3.1 with $f(y)=\|g(y)\|_{2}^{4}$. ∎
Lemma A.2 (Second term $(II)$).
Assume that $\mathbb{E}\left[\|\nabla g(Y)\|_{F}^{2}\right]<\infty$. Then as $\alpha\to 0$,
$\operatorname{Var}\left[\frac{2}{K}\sum_{k=1}^{K}\langle\frac{1}{\sqrt{\alpha}}\omega^{(k)},g(Y+\sqrt{\alpha}\omega^{(k)})\rangle\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}\operatorname{Var}\left[\frac{2}{K}\sum_{k=1}^{K}(\omega^{(k)})^{\intercal}\nabla g(Y)\omega^{(k)}\mid Y\right].$
Proof.
Since all the components of $g$ are almost differentiable, we have
$\langle\frac{1}{\sqrt{\alpha}}\omega,g(Y+\sqrt{\alpha}\omega)\rangle$ $=\frac{1}{\sqrt{\alpha}}\sum_{i=1}^{n}\omega_{i}g_{i}(Y+\sqrt{\alpha}\omega)$
$=\frac{1}{\sqrt{\alpha}}\sum_{i=1}^{n}\omega_{i}\left[g_{i}(Y)+\int_{0}^{1}\sqrt{\alpha}\omega\cdot\nabla g_{i}(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\right]$
$=\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y)+\sum_{i=1}^{n}\omega_{i}\int_{0}^{1}\omega\cdot\nabla g_{i}(Y+t\sqrt{\alpha}\omega)\mathrm{d}t$
$=\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y)+\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\,\omega,$
where $\nabla g$ is the $n\times n$ matrix with the $i$-th row being $\nabla g_{i}$. Averaging over $k=1,\ldots,K$ and using the fact that $\sum_{k=1}^{K}\omega^{(k)}=0$, we have
$\frac{1}{K}\sum_{k=1}^{K}\langle\frac{1}{\sqrt{\alpha}}\omega^{(k)},g(Y+\alpha\omega^{(k)})\rangle=\frac{1}{K}\sum_{k=1}^{K}(\omega^{(k)})^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\,\omega^{(k)}.$
Denote each summand as $\Gamma_{\alpha}(\omega^{(k)},Y)$, where $\Gamma_{\alpha}(\omega,Y)=\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\omega$ and $\Gamma_{0}(\omega,Y)=\omega^{\intercal}\nabla g(Y)\omega$. The claim of the lemma is equivalent to
$\operatorname{Var}\left[\frac{1}{K}\sum_{k=1}^{K}\Gamma_{\alpha}(\omega^{(k)},Y)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\to}}\operatorname{Var}\left[\frac{1}{K}\sum_{k=1}^{K}\Gamma_{0}(\omega^{(k)},Y)\mid Y\right],\text{ as }\alpha\to 0.$
It suffices to show that as $\alpha\to 0$,
$\operatorname{Var}\left[\Gamma_{\alpha}(\omega,Y)\mid Y\right]$ $\stackrel{{\scriptstyle L_{1}}}{{\to}}\operatorname{Var}\left[\Gamma_{0}(\omega,Y)\mid Y\right],$
$\operatorname{Cov}\left[\Gamma_{\alpha}(\omega^{(1)},Y),\Gamma_{\alpha}(\omega^{(2)},Y)\mid Y\right]$ $\stackrel{{\scriptstyle L_{1}}}{{\to}}\operatorname{Cov}\left[\Gamma_{0}(\omega^{(1)},Y),\Gamma_{0}(\omega^{(2)},Y)\mid Y\right].$
The conditional variance and conditional covariance involve the conditional expectation of the following terms:
$\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\,\omega=\sum_{i,j=1}^{n}\int_{0}^{1}\nabla_{j}g_{i}(Y+t\sqrt{\alpha}\omega)\omega_{i}\omega_{j}\mathrm{d}t$
$(\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\omega)^{2}=$
$\qquad\qquad\sum_{i,j,k,l}\int_{[0,1]^{2}}\nabla_{j}g_{i}(Y+t_{1}\sqrt{\alpha}\omega)\nabla_{l}g_{k}(Y+t_{2}\sqrt{\alpha}\omega)\omega_{i}\omega_{j}\omega_{k}\omega_{l}\mathrm{d}t_{1}\mathrm{d}t_{2}$
$({\omega^{(1)}}^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(1)})\mathrm{d}t\omega^{(1)})({\omega^{(2)}}^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega^{(2)})\mathrm{d}t\omega^{(2)})=$
$\qquad\qquad\sum_{i,j,k,l}\int_{[0,1]^{2}}\nabla_{j}g_{i}(Y+t_{1}\sqrt{\alpha}\omega^{(1)})\nabla_{l}g_{k}(Y+t_{2}\sqrt{\alpha}\omega^{(2)})\omega_{i}^{(1)}\omega_{j}^{(1)}\omega_{k}^{(2)}\omega_{l}^{(2)}\mathrm{d}t_{1}\mathrm{d}t_{2}$
Consider the quantities
$\omega_{i}\omega_{j}\int_{0}^{1}\nabla_{j}g_{i}(Y+t\sqrt{\alpha}\omega)\mathrm{d}t$
$\omega_{i}\omega_{j}\omega_{k}\omega_{l}\int_{0}^{1}\int_{0}^{1}\nabla_{j}g_{i}(Y+t_{1}\sqrt{\alpha}\omega)\nabla_{l}g_{k}(Y+t_{2}\sqrt{\alpha}\omega)\mathrm{d}t_{1}\mathrm{d}t_{2}$
$\omega_{i}^{(1)}\omega_{j}^{(1)}\omega_{k}^{(2)}\omega_{l}^{(2)}\int_{0}^{1}\int_{0}^{1}\nabla_{j}g_{i}(Y+t_{1}\sqrt{\alpha}\omega^{(1)})\nabla_{l}g_{k}(Y+t_{2}\sqrt{\alpha}\omega^{(2)})\mathrm{d}t_{1}\mathrm{d}t_{2}.$
We need to show that their conditional expectation $\mathbb{E}\left[\cdot\mid Y\right]$ converges in $L_{1}$ to their corresponding expected value at $\alpha=0$. For the first term, we apply Lemma D.3 with $f=\nabla_{j}g_{i}$ and $h(\omega)=\omega_{i}\omega_{j}$. For the second term, we apply Lemma D.4 with $f_{1}=\nabla_{j}g_{i}$, $f_{2}=\nabla_{l}g_{k}$, and $h(\omega)=\omega_{i}\omega_{j}\omega_{k}\omega_{l}$. For the third term, we can apply Lemma D.4 as well.
A.2 Proof of Theorem 3.4
Proof of Theorem 3.4.
Note that
$\mathbb{E}\left[\mathrm{CV}_{\alpha}\mid Y\right]$ $=\|Y\|_{2}^{2}+\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega)\|_{2}^{2}\mid Y\right]-2Y^{\intercal}\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega)\mid Y\right]$
$+\frac{2}{\sqrt{\alpha}}\mathbb{E}\left[\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right].$
It suffices to show that as $\alpha\to 0$,
$\mathbb{E}\left[\|g(Y+\sqrt{\alpha}\omega)\|_{2}^{2}\mid Y\right]$ $\stackrel{{\scriptstyle L_{2}}}{{\to}}\|g(Y)\|_{2}^{2},$
$\mathbb{E}\left[g(Y+\sqrt{\alpha}\omega)^{\intercal}Y\mid Y\right]$ $\stackrel{{\scriptstyle L_{2}}}{{\to}}g(Y)^{\intercal}Y,$
$\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right]$ $\stackrel{{\scriptstyle L_{2}}}{{\to}}\sigma^{2}\operatorname{tr}(\nabla g(Y)).$
Applying Lemma D.2 with $f(y)=\|g(y)\|_{2}^{4}$ and $h(\omega)=1$ proves the first line. Applying Lemma D.2 with $f(y)=g_{i}(y)g_{j}(y)$ and $h(\omega)=1$ ($1\leq i,j\leq n$) proves the first two lines. For the third limit, recall that
$\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)=\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y)+\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\omega.$
Because $\mathbb{E}\left[\omega^{\intercal}g(Y)\mid Y\right]=0$, we have
$\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right]$ $=\mathbb{E}\left[\omega^{\intercal}\int_{0}^{1}\nabla g(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\omega\mid Y\right]$
$=\sum_{i,j=1}^{n}\mathbb{E}\left[\int_{0}^{1}\nabla_{j}g_{i}(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\omega_{i}\omega_{j}\mid Y\right].$
Applying Lemma D.3 with $f(y)=\nabla_{j}g_{i}(y)\nabla_{l}g_{k}(y)$ and $h(\omega)=\omega_{i}\omega_{j}\omega_{k}\omega_{l}$ ($1\leq i,j,k,l\leq n$) proves the third line. ∎
Appendix B Proofs for Section 4
B.1 Proof of Proposition 4.1
Proof of Proposition 4.1.
Note that
$\widetilde{\mathrm{CV}}_{\alpha}$ $=\|Y-g*\varphi_{\alpha\sigma^{2}}(Y)\|_{2}^{2}+\frac{2}{\sqrt{\alpha}}\mathbb{E}\left[\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right],$
$\mathrm{SURE}(g*\varphi_{\alpha\sigma^{2}})$ $=\|Y-g*\varphi_{\alpha\sigma^{2}}(Y)\|_{2}^{2}+2\sigma^{2}\nabla\cdot(g*\varphi_{\alpha\sigma^{2}})(Y).$
It suffices to show that
$\sigma^{2}\nabla\cdot(g*\varphi_{\alpha\sigma^{2}})(Y)=\frac{1}{\sqrt{\alpha}}\mathbb{E}\left[\omega^{\intercal}g(Y+\sqrt{\alpha}\omega)\mid Y\right].$
This is true because
$\nabla\cdot(g*\varphi_{\alpha\sigma^{2}})(y)$ $=\sum_{i=1}^{n}\nabla_{y_{i}}g_{i}*\varphi_{\alpha\sigma^{2}}(y)=\sum_{i=1}^{n}\nabla_{y_{i}}\int_{\mathbb{R}}g_{i}(x)\varphi_{\alpha\sigma^{2}}(y-x)\mathrm{d}x$
$=\sum_{i=1}^{n}\int_{\mathbb{R}}g_{i}(x)(-\frac{y_{i}-x}{\alpha\sigma^{2}})\varphi_{\alpha\sigma^{2}}(y-x)\mathrm{d}x$
$=-\frac{1}{\alpha\sigma^{2}}\int g(x)^{\intercal}(y-x)\varphi_{\alpha\sigma^{2}}(y-x)\mathrm{d}x$
$=\frac{1}{\alpha\sigma^{2}}\int g(y+\varepsilon)^{\intercal}\varepsilon\varphi_{\alpha\sigma^{2}}(\varepsilon)\mathrm{d}\varepsilon$
$=\frac{1}{\alpha\sigma^{2}}\mathbb{E}_{\varepsilon\sim\mathcal{N}(0,\alpha\sigma^{2}I)}\left[g(y+\varepsilon)^{\intercal}\varepsilon\right]$
$=\frac{1}{\sqrt{\alpha}\sigma^{2}}\mathbb{E}_{\omega\sim\mathcal{N}(0,\sigma^{2}I)}\left[g(y+\sqrt{\alpha}\omega)^{\intercal}\omega\right].$
Appendix C Proofs for Section 5
C.1 Proof of Theorem 5.1
Proof of Theorem 5.1.
The bias can be decomposed as
$|\mathbb{E}\left[\mathrm{CV}_{n,\alpha}\right]-\mathrm{PE}_{n}(g)|$
$\qquad=\Big{|}\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n}+\sqrt{\alpha}\omega)-{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)^{\intercal}(T_{n}-\frac{1}{\sqrt{\alpha}}\omega)-{\mathfrak{A}}_{n}(T_{n})+{\mathfrak{g}}_{n}(T_{n})^{\intercal}m_{n}\right]\Big{|}$
$\qquad\leq\Big{|}\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n}+\sqrt{\alpha}\omega)\right]-\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n})\right]\Big{|}+\Big{|}\mathbb{E}\left[{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)^{\intercal}T_{n}-{\mathfrak{g}}_{n}(T_{n})^{\intercal}T_{n}\right]\Big{|}$
$\qquad\qquad+\Big{|}\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)-{\mathfrak{g}}_{n}(T_{n})^{\intercal}(T_{n}-m_{n})\right]\Big{|}.$ (21)
Because the density $q_{n}$ of $T_{n}$ satisfies Assumption 3, and ${\mathfrak{g}}_{n}$, ${\mathfrak{A}}_{n}$ are integrable w.r.t. $q_{n}$, we can apply Lemma D.2 to show that, for $n\geq N_{0}$,
$\lim_{\alpha\to 0}\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n}+\sqrt{\alpha}\omega)\right]=\mathbb{E}\left[{\mathfrak{A}}_{n}(T_{n})\right],$
$\lim_{\alpha\to 0}\mathbb{E}\left[{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)^{\intercal}T_{n}\right]=\mathbb{E}\left[{\mathfrak{g}}_{n}(T_{n})^{\intercal}T_{n}\right].$
Thus, the first two terms in the upper bound (21) vanish as $\alpha\to 0$. It remains to show that
$\lim_{n\to\infty}\lim_{\alpha\to 0}\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)-{\mathfrak{g}}_{n}(T_{n})^{\intercal}(T_{n}-m_{n})\right]=0.$ (22)
Because ${\mathfrak{g}}_{n}$ is weakly differentiable, we can write
$\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega)\right]=\mathbb{E}\left[\frac{1}{\sqrt{\alpha}}\omega^{\intercal}{\mathfrak{g}}_{n}(T_{n})\right]+\mathbb{E}\left[\omega^{\intercal}\int_{0}^{1}\nabla{\mathfrak{g}}_{n}(T_{n}+t\sqrt{\alpha}\omega)\mathrm{d}t\omega\right].$
The first expectation is 0. For the second term, Lemma D.3 shows that
$\lim_{\alpha\to 0}\mathbb{E}\left[\omega^{\intercal}\int_{0}^{1}\nabla{\mathfrak{g}}_{n}(T_{n}+t\sqrt{\alpha}\omega)\mathrm{d}t\omega\right]=\mathbb{E}\left[\omega^{\intercal}\nabla{\mathfrak{g}}_{n}(T_{n})\omega\right]=\mathbb{E}\left[\nabla\cdot{\mathfrak{g}}_{n}(T_{n})\right].$
By Assumption 2, we have
$\lim_{n\to\infty}\mathbb{E}\left[\nabla\cdot{\mathfrak{g}}_{n}(T_{n})\right]-\mathbb{E}\left[{\mathfrak{g}}_{n}(T_{n})^{\intercal}(T_{n}-m_{n})\right]=0.$
This proves Equation (22). ∎
C.2 Proof of Theorem 5.2
Proof of Theorem 5.2.
We write
$\mathrm{CV}_{n,\alpha}=(I)+(II),$
where
$(I)$ $=\frac{1}{K}\sum_{k=1}^{K}\Big{\{}\mathfrak{A}_{n}(T_{n}+\sqrt{\alpha}\omega^{(k)})-\mathfrak{g}_{n}(T_{n}+\sqrt{\alpha}\omega^{(k)})^{\intercal}T_{n}\Big{\}}-n^{-1/2}\log h_{n}(Y_{n})$
$(II)$ $=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{\sqrt{\alpha}}{\mathfrak{g}}_{n}(T_{n}+\sqrt{\alpha}\omega^{(k)})^{\intercal}\omega^{(k)}$
Because ${\mathfrak{A}}_{n}^{2},\|{\mathfrak{g}}_{n}\|_{2}^{4}$ are integrable under $q_{n}$, and $q_{n}$ satisfies Assumption 3, we can apply Lemma D.2 with $f(y)={\mathfrak{A}}_{n}^{2}(y)$ and $f(y)=\|{\mathfrak{g}}_{n}\|_{2}^{4}$ to show that, when $n\geq N_{0}$,
$\lim_{\alpha\to 0}\mathbb{E}\left[\operatorname{Var}\left[(I)\mid Y_{n}\right]\right]=0.$
For the second term, because ${\mathfrak{g}}_{n}$ is weakly differentiable, we can write
$(II)$ $=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{\sqrt{\alpha}}(\omega^{(k)})^{\intercal}{\mathfrak{g}}_{n}(T_{n})+\frac{1}{K}\sum_{k=1}^{K}\int_{0}^{1}\omega^{(k)}\cdot\nabla{\mathfrak{g}}_{n}(T_{n}+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\omega^{(k)}$
$=\frac{1}{K}\sum_{k=1}^{K}\int_{0}^{1}\omega^{(k)}\cdot\nabla{\mathfrak{g}}_{n}(T_{n}+t\sqrt{\alpha}\omega^{(k)})\mathrm{d}t\omega^{(k)},$
where the second equality is due to the zero-sum constraint $\sum_{k=1}^{K}\omega^{(k)}=0$. The rest of the proof is similar to the proof of Theorem 3.3. ∎
Appendix D Technical lemmas
Lemma D.1.
Let $p$ be a continuous density on $\mathbb{R}^{n}$ with respect to the Lebesgue measure. Suppose there exist a constant $L>0$ such that for all $x,x^{\prime}\in\mathbb{R}^{n}$,
$|\log p(x)-\log p(x^{\prime})|\leq L\|x-x^{\prime}\|_{2}^{2}.$
Let $\varphi$ be the density of the standard normal distribution $\mathcal{N}(0,I_{n})$ with respect to the Lebesgue measure. Then the following results hold:
  1. (i)
    Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be $L_{1}$-integrable with respect to $p$. Let $h:\mathbb{R}^{n}\to\mathbb{R}$ be be integrable w.r.t. $\mathcal{N}(0,(1+\delta_{0})I_{n})$ for some $\delta_{0}>0$. Then there exists $\varepsilon_{0}>0$, such that for all $\varepsilon\in[0,\varepsilon_{0}]$,
    $|h(\omega)|\varphi(\omega)\int_{\mathbb{R}^{n}}|f(x+\varepsilon\omega)|p(x)\mathrm{d}x\leq q(\omega),$
    where $q$ is an $L_{1}$-integrable function on $\mathbb{R}^{n}$.
  2. (ii)
    Let $f_{1},f_{2}$ be $L_{2}$-integrable functions with respect to $p$. Let $h:\mathbb{R}^{n}\to\mathbb{R}$ be integrable w.r.t. $\mathcal{N}(0,(1+\delta_{0})I_{n})$ for some $\delta_{0}>0$. Then there exists $\varepsilon_{0}>0$, such that when $\varepsilon_{1},\varepsilon_{2}\in[0,\varepsilon_{0}]$,
    $|h(\omega)|\varphi(\omega)\int|f_{1}(x+\varepsilon_{1}\omega)f_{2}(x+\varepsilon_{2}\omega)|p(x)\mathrm{d}x\leq q(\omega),$
    where $q$ is an $L_{1}$-integrable function on $\mathbb{R}^{n}$.
In particular, if $p$ is a Gaussian density, then $\log p$ is a quadratic function, so the condition on $\log p$ in the lemma is satisfied.
Proof.
By the assumption on $\log p$, we have
$|\log p(x-\varepsilon\omega)-\log p(x)|\leq L\varepsilon^{2}\|\omega\|_{2}^{2}.$
Then
$\frac{p(x-\varepsilon\omega)}{p(x)}=\exp(\log p(x-\varepsilon\omega)-\log p(x))\leq\exp(L\varepsilon^{2}\|\omega\|_{2}^{2}).$
Thus, we obtain
$\varphi(\omega)\frac{p(x-\varepsilon\omega)}{p(x)}\leq\frac{1}{(2\pi)^{n/2}}\exp(-(\frac{1}{2}-L\varepsilon^{2})\|\omega\|_{2}^{2}).$
Take $\varepsilon_{0}=(\frac{\delta_{0}}{4L(1+\delta_{0})})^{1/2}$. When $\varepsilon\leq\varepsilon_{0}$, we have $\frac{1}{2}-L\varepsilon^{2}\geq\frac{1}{2}-L\varepsilon_{0}^{2}=\frac{1+\delta_{0}/2}{2(1+\delta_{0})}>\frac{1}{2(1+\delta_{0})}$. Define $q(\omega):=|h(\omega)|\frac{1}{(2\pi)^{n/2}}\exp(-\frac{1}{2(1+\delta_{0})}\|\omega\|_{2}^{2})$, which is an integrable function by the assumption. We obtain
$|h(\omega)|\varphi(\omega)p(x-\varepsilon\omega)$ $\leq|h(\omega)|\frac{1}{(2\pi)^{n/2}}\exp(-(\frac{1}{2}-L\varepsilon^{2})\|\omega\|_{2}^{2})\cdot p(x)$
$\leq|h(\omega)|\frac{1}{(2\pi)^{n/2}}\exp(-\frac{1}{2(1+\delta_{0})}\|\omega\|_{2}^{2})\cdot p(x)$
$=q(\omega)p(x).$
Therefore,
$|h(\omega)|\varphi(\omega)\int|f(x+\varepsilon\omega)|p(x)\mathrm{d}x$ $=|h(\omega)|\varphi(\omega)\int|f(x)|p(x-\varepsilon\omega)\mathrm{d}x$
$=\int|f(x)||h(\omega)|\varphi(\omega)p(x-\varepsilon\omega)\mathrm{d}x$
$\leq\int|f(x)|q(\omega)p(x)\mathrm{d}x$
$=q(\omega)\cdot\int|f(x)|p(x)\mathrm{d}x,$
which is integrable because $q$ is integrable and $\int|f(x)|p(x)\mathrm{d}x<\infty$. This proves the first claim.
Next, suppose $f_{1}^{2},f_{2}^{2}$ are integrable functions with respect to $p$. From the first claim, there exists $\varepsilon_{0}$, such that for all $\varepsilon_{1},\varepsilon_{2}\in[0,\varepsilon_{0}]$,
$|h(\omega)|\varphi(\omega)\int f_{1}(x+\varepsilon_{1}\omega)^{2}p(x)\mathrm{d}x\leq q_{1}(\omega),$
$|h(\omega)|\varphi(\omega)\int f_{2}(x+\varepsilon_{2}\omega)^{2}p(x)\mathrm{d}x\leq q_{2}(\omega),$
where $q_{1},q_{2}$ are integrable functions. By the Cauchy-Schwarz inequality, we have
$|h(\omega)|\varphi(\omega)\int|f_{1}(x+\varepsilon_{1}\omega)f_{2}(x+\varepsilon_{2}\omega)|p(x)\mathrm{d}x$
$\qquad\leq\left(|h(\omega)|\varphi(\omega)\int f_{1}(x+\varepsilon_{1}\omega)^{2}p(x)\mathrm{d}x\right)^{1/2}\cdot\left(|h(\omega)|\varphi(\omega)\int f_{2}(x+\varepsilon_{2}\omega)^{2}p(x)\mathrm{d}x\right)^{1/2}$
$\qquad\leq q_{1}(\omega)^{1/2}q_{2}(\omega)^{1/2},$
which is $L_{1}$ integrable, proving the second claim. ∎
Lemma D.2.
Let $p$ be a density on $\mathbb{R}^{n}$ satisfying the condition in Lemma D.1. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function that is integrable w.r.t. the density $p$. Let $h:\mathbb{R}^{n}\to\mathbb{R}$ be a function that is integrable w.r.t. $\mathcal{N}(0,(1+\delta_{0})I_{n})$ for some $\delta_{0}>0$. Then as $\alpha\downarrow 0$,
$\mathbb{E}\left[f(Y+\sqrt{\alpha}\omega)h(\omega)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\rightarrow}}f(Y)\mathbb{E}\left[h(\omega)\right],$
where the expectation is taken over $\omega\sim\mathcal{N}(0,I_{n})$, and the $L_{1}$ convergence is with respect to $Y\sim p$.
Proof.
We start from the inequality
$\Big{|}\int(f(y+\sqrt{\alpha}\omega)h(\omega)-f(y)h(\omega))\varphi(\omega)\mathrm{d}\omega\Big{|}$ $\leq\int|f(y+\sqrt{\alpha}\omega)-f(y)|\cdot|h(\omega)|\varphi(\omega)\mathrm{d}\omega.$
Multiplying by $p(y)$, integrating over $y$, and applying Fubini’s theorem, we have
$\int\Big{|}\int(f(y+\sqrt{\alpha}\omega)h(\omega)-f(y)h(\omega))\varphi(\omega)\mathrm{d}\omega\Big{|}p(y)\mathrm{d}y$
$\qquad\leq\int\int|f(y+\sqrt{\alpha}\omega)-f(y)|\cdot|h(\omega)|\varphi(\omega)\mathrm{d}\omega p(y)\mathrm{d}y$
$\qquad=\int|h(\omega)|\varphi(\omega)\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y\mathrm{d}\omega.$
By Lemma D.1, there exists $\varepsilon_{0}>0$ such that, for all $\varepsilon\in[0,\varepsilon_{0}]$,
$|h(\omega)|\varphi(\omega)\int|f(y+\varepsilon\omega)|p(y)\mathrm{d}y\leq\bar{f}(\omega),$
where $\bar{f}(\omega)$ is $L_{1}$ integrable. When $\sqrt{\alpha}\leq\varepsilon_{0}$, we have
$|h(\omega)|\varphi(\omega)\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y$ $\leq 2\bar{f}(\omega).$
Thus, $|h(\omega)|\varphi(\omega)\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y$ is dominated by an integrable function. Moreover, $\lim_{\alpha\to 0}\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y=0$ (Lebesgue differentiation theorem or mean continuity theorem; proof by the fact that compactly supported continuous functions are dense in $L_{1}$). Applying the dominated convergence theorem, we have
$\lim_{\alpha\to 0}\int|h(\omega)|\varphi(\omega)\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y\mathrm{d}\omega$
$\quad=\int|h(\omega)|\varphi(\omega)\lim_{\alpha\to 0}\int|f(y+\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y\mathrm{d}\omega$
$\quad=0.$
Lemma D.3.
Under the same condition as in Lemma D.2, as $\alpha\downarrow 0$, we have
$\mathbb{E}\left[\int_{0}^{1}f(Y+t\sqrt{\alpha}\omega)\mathrm{d}t\cdot h(\omega)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\rightarrow}}f(Y)\mathbb{E}\left[h(\omega)\right],$
where the expectation is taken over $\omega\sim\mathcal{N}(0,I_{n})$, and the $L_{1}$ convergence is with respect to $Y\sim p$.
Proof.
We start from
$\Big{|}\int_{0}^{1}\int(f(y+t\sqrt{\alpha}\omega)h(\omega)-f(y)h(\omega))\varphi(\omega)\mathrm{d}\omega\mathrm{d}t\Big{|}$
$\qquad\leq\int_{0}^{1}\int|f(y+t\sqrt{\alpha}\omega)-f(y)|\cdot|h(\omega)|\varphi(\omega)\mathrm{d}\omega\mathrm{d}t.$
Integrating over $y$ and applying Fubini’s theorem, we have
$\int\Big{|}\int_{0}^{1}\int(f(y+t\sqrt{\alpha}\omega)h(\omega)-f(y)h(\omega))\varphi(\omega)\mathrm{d}\omega\mathrm{d}t\Big{|}p(y)\mathrm{d}y$
$\qquad\leq\int_{0}^{1}\int|h(\omega)|\varphi(\omega)\int|f(y+t\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y\mathrm{d}\omega\mathrm{d}t.$
When $\sqrt{\alpha}<\varepsilon_{0}$, $t\sqrt{\alpha}<\varepsilon_{0}$ for all $t\in[0,1]$. Following a similar argument as in the proof of Lemma D.2, we have $|h(\omega)|\varphi(\omega)\int|f(y+t\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y\leq 2\bar{f}(\omega)$, which is integrable. Moreover, $\lim_{\alpha\to 0}\int|f(y+t\sqrt{\alpha}\omega)-f(y)|p(y)\mathrm{d}y=0$ (mean continuity theorem). By dominated convergence theorem, the last display converges to 0 as $\alpha\to 0$. ∎
Lemma D.4.
Suppose $p$ is a density satisfying the condition in Lemma D.1. Suppose $f_{1}(y)^{2},f_{2}(y)^{2}$ are integrable functions w.r.t. $p$. Suppose $h$ is integrable w.r.t. $\mathcal{N}(0,(1+\delta_{0})I_{n})$ for some $\delta_{0}>0$. Then as $\alpha\downarrow 0$, we have
$\mathbb{E}\left[\int_{0}^{1}\int_{0}^{1}f_{1}(Y+t_{1}\sqrt{\alpha}\omega)f_{2}(Y+t_{2}\sqrt{\alpha}\omega)\mathrm{d}t_{1}\mathrm{d}t_{2}\cdot h(\omega)\mid Y\right]\stackrel{{\scriptstyle L_{1}}}{{\rightarrow}}f_{1}(Y)f_{2}(Y)\mathbb{E}\left[h(\omega)\right],$
where the expectation is taken over $\omega\sim\mathcal{N}(0,I_{n})$, and the $L_{1}$ convergence is with respect to $Y\sim p$.
Proof.
Because $f_{1}^{2},f_{2}^{2}$ are integrable w.r.t. $p$, applying the second claim in Lemma D.1, there exists $\varepsilon_{0}$ such that, when $\varepsilon_{1},\varepsilon_{2}\in[0,\varepsilon_{0}]$,
$|h(\omega)|\varphi(\omega)\int|f_{1}(y+\varepsilon_{1}\omega)f_{2}(y+\varepsilon_{2}\omega)|p(y)\mathrm{d}y\leq\bar{f}(\omega),$
where $\bar{f}(\omega)$ is integrable. Following the same argument as in the proof of Lemma D.3, we have
$\int\bigg{|}\int_{0}^{1}\int_{0}^{1}\int(f_{1}(y+t_{1}\sqrt{\alpha}\omega)f_{2}(y+t_{2}\sqrt{\alpha}\omega)h(\omega)-f_{1}(y)f_{2}(y)h(\omega))\varphi(\omega)\mathrm{d}\omega\mathrm{d}t_{1}\mathrm{d}t_{2}\bigg{|}p(y)\mathrm{d}y$
$\quad\leq\int_{0}^{1}\int_{0}^{1}\int|h(\omega)|\varphi(\omega)\int\big{|}f_{1}(y+t_{1}\sqrt{\alpha}\omega)f_{2}(y+t_{2}\sqrt{\alpha}\omega)-f_{1}(y)f_{2}(y)\big{|}p(y)\mathrm{d}y\mathrm{d}\omega\mathrm{d}t_{1}\mathrm{d}t_{2}.$
Following a similar argument as in the proof of Lemma D.3, when $\sqrt{\alpha}<\varepsilon_{0}$, the function
$|h(\omega)|\varphi(\omega)\int|f_{1}(y+t_{1}\sqrt{\alpha}\omega)f_{2}(y+t_{2}\sqrt{\alpha}\omega)-f_{1}(y)f_{2}(y)|p(y)\mathrm{d}y$
is dominated by an integrable function. Since $\lim_{\alpha\to 0}\int|f_{1}(y+t_{1}\sqrt{\alpha}\omega)f_{2}(y+t_{2}\sqrt{\alpha}\omega)-f_{1}(y)f_{2}(y)|p(y)\mathrm{d}y=0$ (mean continuity theorem), the proof is completed by applying the dominated convergence theorem. ∎
Lemma D.5.
Suppose $x,y\sim\mathcal{N}(0,I_{n})$ and $\operatorname{Cov}\left[x,y\right]=\rho I_{n}$. For a matrix $A$, we have
$\operatorname{Cov}\left[x^{\intercal}Ax,y^{\intercal}Ay\right]=\rho^{2}\operatorname{Var}\left[x^{\intercal}Ax\right]=\rho^{2}(\|A\|_{F}^{2}+\operatorname{tr}(A^{2})).$
Proof.
Let $x,z\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(0,I_{n})$ and $y=\rho x+\widetilde{\rho}z$ where $\widetilde{\rho}=\sqrt{1-\rho^{2}}$. Then $y\sim\mathcal{N}(0,I_{n})$ and $\operatorname{Cov}\left[x,y\right]=\rho I_{n}$. We have
$\mathbb{E}\left[(x^{\intercal}Ax)(y^{\intercal}Ay)\right]$ $=\mathbb{E}\left[(x^{\intercal}Ax)(\rho^{2}x^{\intercal}Ax+\widetilde{\rho}^{2}z^{\intercal}Az)\right]$
$=\rho^{2}\mathbb{E}\left[(x^{\intercal}Ax)^{2}\right]+\widetilde{\rho}^{2}\mathbb{E}\left[(x^{\intercal}Ax)\right]\mathbb{E}\left[(z^{\intercal}Az)\right]$
$=\rho^{2}\mathbb{E}\left[(x^{\intercal}Ax)^{2}\right]+\widetilde{\rho}^{2}\operatorname{tr}(A)^{2}.$
Note that
$\mathbb{E}\left[(x^{\intercal}Ax)^{2}\right]$ $=\mathbb{E}\left[\sum_{i,j,k,l}A_{ij}A_{kl}x_{i}x_{j}x_{k}x_{l}\right]$
$=\mathbb{E}\left[\sum_{i}A_{ii}^{2}x_{i}^{4}+\sum_{i\neq j}A_{ij}^{2}x_{i}^{2}x_{j}^{2}+\sum_{i\neq j}A_{ij}A_{ji}x_{i}^{2}x_{j}^{2}+\sum_{i\neq j}A_{ii}A_{jj}x_{i}^{2}x_{j}^{2}\right]$
$=3\sum_{i}A_{ii}^{2}+\sum_{i\neq j}A_{ij}^{2}+\sum_{i\neq j}A_{ij}A_{ji}+\sum_{i\neq j}A_{ii}A_{jj}$
$=\sum_{i,j}A_{ij}^{2}+\sum_{i,j}A_{ij}A_{ji}+\sum_{i,j}A_{ii}A_{jj}$
$=\operatorname{tr}(A^{\intercal}A)+\operatorname{tr}(AA)+\operatorname{tr}(A)^{2}.$
Moreover, $\mathbb{E}\left[x^{\intercal}Ax\right]=\operatorname{tr}(A)$. So we have
$\operatorname{Cov}\left[x^{\intercal}Ax,y^{\intercal}Ay\right]$ $=\rho^{2}\mathbb{E}\left[(x^{\intercal}Ax)^{2}\right]+\widetilde{\rho}^{2}\operatorname{tr}(A)^{2}-\operatorname{tr}(A)^{2}$
$=\rho^{2}(\operatorname{tr}(A^{\intercal}A)+\operatorname{tr}(AA)+\operatorname{tr}(A)^{2})+\widetilde{\rho}^{2}\operatorname{tr}(A)^{2}-\operatorname{tr}(A)^{2}$
$=\rho^{2}(\operatorname{tr}(A^{\intercal}A)+\operatorname{tr}(AA))$
$=\rho^{2}\operatorname{Var}\left[x^{\intercal}Ax\right].$